### v1

# Graph Neural Network Enhanced Retrieval for Question Answering of LLMs

Abstract
[…]
1 Introduction
Large language models (LLMs) continue to struggle with factual errors when encountering knowledge intensive questions [7, 19, 9]. Although retrieval-augmented LLMs [16, 15] have improved factuality and precision of question answering by including relevant passages, there remains a persistent challenge in accurately capturing all the relevant supporting passages when encountering complex knowledge-intensive questions. This limitation can be attributed to the inherent information asymmetry in complex questions. In particular, the questions tend to consist of elaborate background details, leaving only a small portion dedicated to specific inquiries. As an example, in question: ‘Why did crime rise on Mars after the Mafia’s arrival?’, the majority part delves into the background (i.e., ‘crime rise …… Mafia’s arrival’) with only a few words requesting the reason (i.e., ‘Why’), which consequently retrieves passages that focus on the details of crime rising instead of the reason for it.
This phenomenon also frequently arises in multi-hop reasoning questions. Considering the sample question ‘Where was the performer of song Left & Right (Song) born?’ in Fig. 1, it becomes apparent that while we can retrieve the knowledge that the performer of song Left & Right is D’Angelo, his birthplace remains absent. Previous works [30, 23, 35] have attempted to address this issue by incorporating multi-hop reasoning or question rewriting into retrieval processes, which enables them to retrieve information based on prior reasoning outcomes. LLMs, however, often generate plausible reasons and incorrect rewriting questions without accurate prior domain knowledge of the given question [7, 36], thus affecting the subsequent retrieval process.
Figure 1: Overview of comparison between Dense retrieval and GNN-Ret. The shared keywords and ground-truth answers are highlighted in yellow and green, respectively. By considering the relatedness between passages, GNN-Ret can retrieve all the supporting passages for QA.
One reason why existing methods struggle to handle the information asymmetry is their tendency to consider passages in isolation [14] and retrieve the supporting passages based on semantic distances, making it difficult to retrieve all the supporting passages, especially those containing only few words for inquiry. However, the supporting passages of background and inquiry information are usually correlated. They can be structure-related when located in the same section or document, e.g., the happened event (i.e., crime rises on Mars) and its corresponding reason are located in the same section but different passages. Also, they can be keyword-related by sharing the same keywords or entities, e.g., the passages of background information (performer of song Left &\&& Right) and inquiry information (birthplace of D’Angelo) in Fig. 1 share the same keyword: D’Angelo. By considering the relatedness between passages, it is possible to retrieve all the supporting passages even when they have significant semantic differences from the query. In this work, we aim to enhance retrieval by taking the relatedness between passages into account.
Contributions. To establish the relatedness between passages for retrieval purposes, this study initially constructs a graph of passages by connecting individual passages based on both structural information and shared keywords, with each passage as a node in this graph. The key challenge lies in how to effectively leverage the passage of graphs to enhance retrieval coverage. Graph neural networks (GNNs) are neural networks tailored to analyze graph data and adeptly grasp the relationships between nodes and edges [26]. Thus, we propose to leverage a GNN to enhance the retrieval process by effectively capturing the relatedness between passages and name this method GNN-Ret. The GNN facilitates the integration of semantic distances between related passages, thus enabling the semantic distances of passages containing background information to impact the retrieval of passages relevant to the inquiry.
Figure 2: Accuracy and average query times of LLMs for our proposed methods and baselines on 2WikiMQA dataset.
To address multi-hop reasoning questions, we further propose a retrieval method, named RGNN-Ret, which leverages a recurrent graph neural network (RGNN) to enhance the retrieval process at each step by integrating the retrievals from previous steps through the interconnected graphs of passages. This boosts the retrieval coverage for supporting passages over steps even when LLMs generate incorrect reasons or subquestions for retrieval.
Through the experiments on four benchmark datasets, we demonstrate the effectiveness of GNN in enhancing retrieval coverage for supporting facts and thus improving accuracy for QA tasks. For example, as shown in Fig. 2, our proposed GNN-Ret with a single query of LLMs significantly outperforms baselines in accuracy on the 2WikiMQA dataset [6], including those methods that require multiple queries of LLMs, which demonstrates the effectiveness of GNNs in improving retrieval coverage of supporting passages. Moreover, by extending GNN to multi-hop reasoning questions, our proposed RGNN-Ret achieves state-of-the-art accuracy.
2 Related Works
Retrieval-augmented LLM. QA [31] is a task that often requires external and up-to-date knowledge sources to answer factoid-based questions. Retrieve-and-read is the basic framework for these questions [5, 37]. For retrieval, sparse retrievers, e.g., TF-IDF and BM25 [24], or dense retrievers, e.g., DPR [13] and Contriever [8], are used to compute lexical distances or semantic distances between passages and the question. The passages with smaller distances are retrieved, which are then prefixed with the question for factual answering [16]. However, this framework treats passages as isolated units, making it difficult to retrieve all the supporting passages for the question in a single step. To address this limitation, many approaches have been proposed to integrate the retrieval and reasoning processes to improve the retrieval of supporting passages [5, 27, 3, 30, 11, 35, 23]. They first prompt LLMs to generate the next-step reason or subquestion and then use it to guide retrieval or knowledge search for QA. However, a challenge in these methodologies pertains to the potential generation of hallucinated reasons or erroneous subquestions by LLMs without accurate domain knowledge of this question, leading to degraded subsequent retrievals [7].
Graph-enhanced LLM. As structured, explicit, and editable representations of knowledge, knowledge graphs (KGs) have been effectively used for boosting retrieval coverage [20] and enhancing reasoning capabilities of LLMs [17, 34, 1, 32, 22, 2]. Instead of naively combining KGs with LLMs, many works directly prompt LLMs to perform reasoning on KGs [28, 10, 18]. They first identified the entities on the knowledge graph and then search reasoning paths accordingly [28, 18, 10]. However, KG-based methods encounter a limitation that they can only handle questions that are effectively represented as KGs. They may struggle with questions that necessitate a comprehensive understanding of lengthy contextual information, such as the mentioned example: the cause of crime increasing. Rather than KGs, a recent study [12] attempts to construct a graph of passages and perform reasoning on this graph. It connects documents according to some specific words and performs reasoning on this graph for QA. In contrast, we utilize the graph of passages to enhance the retrieval process. Another related work, KGP [33], performs retrieval on the graph and search related passages. However, this approach does not consider the integration between related passages and may not effectively enhance retrieval coverage for the supporting passages with poor semantic similarity.
3 Graph Neural Network Enhanced Retrieval for QA
In this section, we aim to enhance the retrieval coverage of supporting passages by leveraging the relatedness between passages for QA tasks. Specifically, we construct a graph of passages by connecting passages that are structural-related and keyword-related, where each node represents a passage (Section 3.2). To exploit the relatedness between passages to enhance retrieval, we utilize a GNN to enable integration between related passages (Section 3.3). For multi-hop questions, we further leverage an RGNN to establish relationship between graphs over steps to enhance retrieval for the multi-hop answering process (Section 3.4).
3.1 Problem Description
The QA system typically adapts a retrieve-and-read routine [16]: the retriever computes semantic distances $D_{q}$ between passages and question $q$ in the embedding space, finds out the relevant passages $P_{q}$ from the whole corpus $P$, and then the reader (LLM) generates the answer $a$ based on them. This process can be formulated as follows:
$D_{q}=\{\phi(q,p)|p\in P\},\quad P_{q}=\text{Retrieve}(q,D_{q}),\quad a=\text{Read}(q,P_{q}),$ (1)
where $\phi(q,p)$ measures the semantic distance between question $q$ and passage $p$. A smaller value of $\phi(q,p)$ indicates higher similarity between $q$ and $p$. Considering the information asymmetry of complex questions, directly using the initial semantic distances of passages $D_{p}$ for retrieval is difficult to retrieve all the supporting passages. Therefore, we propose to process $D_{q}$ by taking the relatedness between passages into account and thus enhance the retrieval for QA.
3.2 Graph of Passages
Establish relationships between related passages. Existing retrieval methods perform retrieval merely relying on semantic distances between passages and the question [5, 37], which may fail to retrieve some supporting passages with large semantic distances to the question due to the information asymmetry. However, these supporting passages are usually related to those with small semantic distances to the question. They can be structural-related, e.g., the happened event (‘crime rising’) and its corresponding reasons are located in the same section. As shown in Fig. 1, they can also be keyword-related, e.g., the supporting passages for background information (‘performer of song Left &\&& Right’) and inquiry information (‘the birthplace of D’Angelo’) share the same keyword (‘D’Angelo’). Therefore, we establish relationship between passages using structural information and shared keywords to construct the graph of passages. Specifically, when chunking the documents, we record the order of passages and connect passages that are physically next to each other in the documents. In addition, we extract keywords from passages by prompting LLMs and connect passages containing the same keywords [20, 33]. As such, the related passages are connected, having the potential to facilitate retrieval.
3.3 GNN Enhanced Retrieval on Graph of Passages
After establishing connections between related passages, we obtain a graph of passages, with each passage as a node. When conducting the retrieval process, we first compute the semantic distances between passages and the question $q$ [5, 37]. The semantic distances of related passages are connected according to the graph of passages. To take relatedness between passages into account, we utilize a GNN to process semantic distances based on the graph of passages and obtain the integrated semantic distances, which benefits the retrieval of supporting passages.
Graph neural network. A graph is defined as an ordered pair $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{G}$ is the set of nodes $v_{i}$ and $\mathcal{E}$ is the set of edges $e_{ij}$ connecting two nodes $v_{i}$ and $v_{j}$. We define $h_{i}^{l}$ as the hidden state (integrated semantic distance) of the $i-$th node for $i\in\{1,\cdots,|\mathcal{V}|\}$ at layer $l$. We input semantic distances of passages into GNN as $h_{i}^{0}$. The set of nodes $\mathcal{N}_{v_{i}}$ stands for the neighbors of node $v_{i}$. For an $L-$layer GNN [26], the hidden state of each node $v_{i}\in\mathcal{V}$ is updated through iterative interactions with neighbors. Specifically, at layer $l$, the received messages $m_{i}^{l}$ of node $v_{i}$ are calculated using the hidden states of its neighbors. Then the received messages $m_{i}^{l}$ and the previous-layer hidden state $h_{i}^{l-1}$ are utilized to compute the hidden state $h_{i}^{l}$ at layer $l$, which are formulated as follows:
$m_{i}^{l}=f_{m}^{l}(h_{j}^{l-1}|\forall j\in\{1,\cdots,|\mathcal{N}_{v_{i}}|\}),\quad h_{i}^{l}=f_{h}^{l}(m_{i}^{l},h_{i}^{l-1}),$ (2)
where functions $f_{m}^{l}(\cdot)$ and $f_{h}^{l}$ aggregate neighbors’ information and update the hidden state, respectively. By processing hidden states for all the passages layer by layer, we obtain the integrated semantic distance $h_{i}^{L}$ for each node $v_{i}\in\mathcal{V}$. We elaborately introduce the GNN for retrieval as below.
Minimum semantic distance as message. Since each passage maintains and shares many keywords with other passages, each passage has a large number of neighbors. Some of them are not relevant to the question, contributing to distracting or misleading information. Therefore, each node $v_{i}$ only receives the message from the neighbor that has the minimum (integrated) semantic distance to the question at layer $l$. The received message of node $v_{i}$ is thus formulated as $m_{i}^{l}=\min_{j\in\mathcal{N}_{v_{i}}}h_{j}^{l-1}$.
Relevant nodes sampling. Given the large scale of the graph constructed by plenty of passages, it is computationally inefficient to propagate information across all nodes. Besides, allowing irrelevant nodes to pass messages to their neighbors may affect the retrieval process. Therefore, we only sample the relevant node set $\mathcal{T}_{K}^{l}$ with the top $K$ smallest (integrated) semantic distances at layer $l$. Only neighbors of relevant nodes receive messages from them and update hidden states. We define $\mathcal{S}_{K}^{l}$ as the neighbor set of the relevant node set $\mathcal{T}_{K}^{l}$. For each node $v_{i}\in\mathcal{S}_{K}^{l}$, the message passed from neighbors that are relevant nodes is reformulated as follows:
$m_{i}^{l}=\min_{j\in\mathcal{N}_{v_{i}},v_{j}\in\mathcal{T}_{K}^{l}}h_{j}^{l-1}.$ (3)
By integrating the message from neighbors, the hidden state of each node $v_{i}\in\mathcal{S}_{K}^{l}$ at layer $l$ is computed using the parameter $\alpha^{l}$ as follows:
$h_{i}^{l}=\alpha^{l}h_{i}^{l-1}+(1-\alpha^{l})m_{i}^{l}.$ (4)
The nodes without received message maintain their hidden states of the previous layer. For the GNN with $L$ layers, we compute the hidden state layer by layer and then obtain the integrated semantic distance $h_{i}^{L}$ for each node $v_{i}$. With GNN, the supporting passages for inquiry information integrate with the small semantic distance from those for background information and thereby obtain smaller integrated semantic distances, promoting them to be retrieved.
Figure 3: Illustration of GNN-Ret and RGNN-Ret.
Hinge objective for GNN. To retrieve all the supporting passages, we aim to reduce their integrated semantic distances more than those of non-target nodes. With the ground-truth set of supporting passages $\mathcal{S}_{Y}$, we first define the average integrated semantic distances of supporting passages and non-target nodes as $\bar{d}_{Y}^{L}=\frac{1}{|\mathcal{S}_{Y}|}\sum_{j\in\mathcal{S}_{Y}}h_{j}^{L}$ and $\bar{d}_{O/Y}^{L}=\frac{1}{|\mathcal{S}_{O/Y}|}\sum_{o\in\mathcal{S}_{O/Y}}h_{o}^{L}$, respectively, where $S_{O}$ is the competitive node set with the top $O$ smallest semantic distances, and $S_{O/Y}$ is the non-target node set by removing the target node set $S_{Y}$ from $S_{O}$. We only consider the competitive node set since the average integrated semantic distance of all the nodes is significantly large and loses effectiveness in training. With the output average semantic distances of supporting passages and non-target nodes, we formulate the hinge objective function with threshold $r$ as follows:
$\ell=\max(0,r+\bar{d}_{Y}^{L}-\bar{d}_{O/Y}^{L}).$ (5)
We update the GNN using the gradient descent, and training details are elaborated in Appendix A.1.
3.4 RGNN Enhanced Retrieval for Multi-hop Reasoning Questions
The information asymmetry phenomenon also frequently arises in complex multi-hop reasoning questions, as analyzed in the Introduction section. Although recent works have incorporated multi-hop reasoning or question rewriting into retrieval [35, 30, 27], LLMs may generate hallucinated reasons or incorrect subquestions in the absence of prior knowledge in this domain [7]. These hallucinated reasons or incorrect subquestions fail to retrieve supporting passages when only considering the retrieval process in isolation for each step. Additionally, they also encounter another challenge: whether to continue the next-step answering process or output the final answer. To overcome these challenges, we first propose a self-critique technique by prompting LLMs to determine the termination of the answering process. To enhance retrieval of supporting passages over steps, we further utilize an RGNN to enable the integration between graphs of passages in different steps.
Self-critique of LLMs. To answer multi-hop reasoning questions, self-critique involves prompting the LLM to generate subquestions and answer them using retrieved passages in a step-by-step manner. As depicted in Fig. 3, at each step, after the LLM generates a subanswer to the subquestion, the question and all the generated subanswers are input into the LLM and critiqued to determine whether the generated subanswers are evident to generate the final answer to the question. If the output is ‘Yes’, the LLM will be prompted to generate the final answer. Otherwise, it will be requested to generate the next-step subquestion. Compared with the prior work, SelfAsk [23], we focus on whether existing evidences is satisfactory for answering the question instead of whether the next-step subquestion is needed. This self-critique technique decomposes the question over steps and enables a more accurate answer to the question.
Recurrent graph neural network. To enhance retrievals over steps, as shown in Fig. 3, we utilize an RGNN to establish relationships between the semantic distances of subquestions. Specifically, with the subquestion $q^{t}$ at step $t$, we first compute $h_{i}^{L,t}$ for each node $v_{i}\in\mathcal{V}$ using the aforementioned $L$-layer GNN with Eq. (3) and (4). To consider the effect of semantic distances from previous subquestions, we integrate $h_{i}^{L,t}$ with the previous-step integrated semantic distance to compute the integrated semantic distance $\hat{h}_{i}^{L,t}$ at step $t$ using the parameter $\beta$. This can be formulated as follows:
$\hat{h}_{i}^{L,t}=\beta h_{i}^{L,t}+(1-\beta)\hat{h}_{i}^{L,t-1},\quad(t>0),$ (6)
where $\hat{h}_{i}^{L,t-1}$ is the integrated semantic distance from the previous step $t-1$ and $\hat{h}_{i}^{L,0}=h_{i}^{L,0}$. The integrated semantic distance $\hat{h}_{i}^{L,t}$ will be used to integrate with ${h}_{i}^{L,t+1}$ of node $v_{i}$ at the next step $t+1$ recurrently until the last step of the answering process $T$. GNN facilitates the integration of semantic distances between related passages at each step, and the recurrent parameter $\beta$ enables the integration of semantic distances of different subquestions across steps, which effectively mitigates the impacts of incorrect subquestions and enhances the retrieval of supporting passages for them.
Hinge objective for RGNN. We adopt a hinge objective for RGNN by considering all the $T$ answering steps for question $q$. With the subquestion $q^{t}$ at step t, the primary objective of RGNN is to reduce the semantic distance of its corresponding supporting passage and make it retrieved. Furthermore, to enhance the retrieval process for subsequent answering steps, we also reduce the semantic distances of supporting passages that should appear at subsequent steps to make them retrieved subsequently. Specifically, at step $t$, we adopt a hinge objective that encourages the average integrated semantic distance of supporting passages that should appear at the current and subsequent steps $t+=\{t,\cdots,T\}$ to be lower than that of the non-target node set. We first define the average integrated semantic distances of supporting passages and non-target nodes at step $t$ as $\bar{d}_{Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{Y^{t+}}|}\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\hat{h}_{j}^{L,t}$ and $\bar{d}_{O/Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{O/Y^{t+}}|}\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\hat{h}_{o}^{L,t}$, respectively, where $\mathcal{S}_{O}^{t}$ is the competitive node set with top $O$ smallest semantic distances, and $\mathcal{S}_{O/Y^{t+}}^{t}$ is the non-target node set by removing the target node set $\mathcal{S}_{Y^{t+}}^{t}$ from $\mathcal{S}_{O}^{t}$. The hinge objective function with threshold $r$ for the RGNN is formulated as follows:
$\ell$ $=\frac{1}{T}\sum_{t\in[T]}\ell^{t}=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (7)
By employing this hinge objective, the RGNN takes into account the retrieval of subsequent answering steps and proactively reduces the semantic distances associated with them. These small semantic distances are then transferred to the next subquestion and enhance the retrieval of the corresponding supporting passages. We update the RGNN (parameters $\alpha^{l}$ and $\beta$) using the gradient descend, and more training details are elaborated in Appendix A.2.
4 Experiments
In this section, we aim to conduct experiments to demonstrate the effectiveness of GNN-Ret and RGNN-Ret as follows: 1) Our proposed GNN-Ret and RGNN-Ret achieve superior accuracy on various QA datasets (Table 1). 2) We conduct ablation studies to select hyperparameters $K$ and $O$ for training GNN and RGNN (Fig. 4) and verify the effectiveness of each component in GNN-Ret and RGNN-Ret (Table 2). 3) GNN-Ret improves the retrieval coverage of supporting passages (Fig. 5).
4.1 Experimental Setups
Evaluation datasets. We measure all the methods on four different QA datasets, including multi-hop Wikipedia reasoning datasets: (1) MuSiQue [29], (2) IIRC [4], (3) 2WikiMQA [6], and a single-hop multi-choice QA dataset: (4) Quality [21]. For each multi-hop dataset, we randomly sample 500 questions from the development set, which provides the ground-truth supporting passages. We use 20 of them to train the GNN and select the rest as test data. We use all the Wikipedia documents of these 500 questions as the retrieval corpus, which is significantly larger than that provided by datasets [^1] Prior works often use 10–20 paragraphs or documents as the retrieval corpus for each question [27, 33].. For each document or article, we split it into multiple passages with a maximum token length of 200 [25]. As evaluation metrics, we calculate the F1 score, exact match (EM) and accuracy (Acc) for multi-hop reasoning datasets. We use the ChatGPT (gpt-3.5-turbo-2023-06-01-preview) to evaluate if the prediction matches with the gold answer. For Quality, we randomly sample 30 articles with overall corresponding 561 questions and use these articles as the retrieval corpus. Since Quality is not a multi-hop dataset, we only validate the accuracy performance of GNN-Ret on it.
Baselines. We compare GNN-Ret and RGNN-Ret with the following baselines: (1) Direct answers questions without retrieved passages. (2) Dense [14] answers questions with retrieved passages using dense retrieval. (3) SelfAsk [23] prompts LLMs to generate the follow-up question and answers it with retrieved passages until generating the final answer. (4) ITER-RETGEN [27] iteratively answers questions with retrieved passages and uses the generated answer for the next-step retrieval until the generation of final answer. (5) IRCoT [30] iteratively prompts LLMs to generate chains of thoughts with retrieved passages and retrieves with the generated reasons until reaching the maximum token number. All the retrieved passages are then used to generate the final answer. (6) KGP [33] first searches seed passages using the question and then prompts LLMs to generate the needed evidence to retrieve the other relevant passages among neighbors of seed nodes using semantic distances.
Implementation details. We use the ChatGPT (gpt-3.5-turbo-2023-06-01-preview) as the LLM backbone for experiments and adopt a temperature of 0 to remove the effect of random sampling. We select mpnet-base-cos-v1222https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1 as the embedding model to compute the semantic distance $=1-\text{cosine similarity}$ following [25] for all the methods. We set up a maximum token number of 3500 for retrieval to leave the space to instruction and demonstrations, and all the methods retrieve the semantically similar passages until reaching the maximum token number. We adopt a one-layer GNN for our proposed methods GNN-Ret and RGNN-Ret. We set up $\alpha^{1}=0.5$ for GNN-Ret and $\alpha^{1}=0.5$ and $\beta=0.9$ for RGNN-Ret when there is no training. When training the RGNN, we set up $r=0.01$, $K=5$, and $O=25$ for GNN-Ret (w. train) and $r=0.01$, $K=5$, and $O=10$ for RGNN-Ret (w. train), respectively, according to the ablation studies in Fig. 4. We do not train the parameters for the Quality dataset without the ground-truth labels of supporting passages. Since training the RGNN requires ground-truth subquestions, we manually generate subquestions for 5 questions sampled from the preset training data. More implementation details are present in Appendix B.
Methods MuSiQue IIRC 2WikiMQA Quality
F1 EM Acc F1 EM Acc F1 EM Acc F1 EM Acc
No Retrieval
Direct 14.1 4.0 9.6 15.8 9.6 16.0 24.4 17.3 24.8 - - 39.5
One-hop Retrieval
Dense 28.6 13.5 23.8 31.6 17.5 40.4 43.4 31.0 42.3 - - 53.4
GNN-Ret (K=5) 31.6 16.5 27.1 32.2 17.6 44.0 47.7 32.7 44.8 - - 55.5
GNN-Ret (K=10) 30.0 14.4 24.4 35.2 19.6 44.0 47.7 32.9 46.3 - - 58.2
GNN-Ret (w. train) 31.1 17.1 27.3 36.1 21.9 44.4 48.1 33.5 46.9 - - -
Multi-hop Retrieval
SelfAsk [23] 28.0 18.3 22.9 37.1 29.6 36.9 51.2 41.5 44.8 - - -
ITER-RETGEN [27] 30.0 16.7 27.1 32.6 20.0 39.6 45.6 33.3 43.8 - - -
IRCoT [30] 29.9 12.5 27.5 32.5 18.8 41.0 46.0 30.8 44.6 - - -
KGP [33] 30.0 15.4 24.0 33.9 18.5 42.1 45.5 31.9 45.2 - - -
RGNN-Ret 32.9 20.8 31.3 43.4 28.3 48.1 59.7 43.3 55.8 - - -
RGNN-Ret (w. train) 34.8 21.9 31.3 42.8 27.6 48.4 61.4 45.2 55.6 - - -
Table 1: F1 / EM / Acc for different QA methods with ChatGPT on four QA datasets. The best and second best scores are highlighted in bold and underline, respectively
4.2 Main Results
Table 1 summarizes the results for our proposed methods and baselines on four QA datasets using ChatGPT. The results show that our proposed GNN-Ret significantly outperforms baselines Dense and Direct in terms of F1, EM, and accuracy on all these four QA datasets with a single retrieval. For instance, compared with Dense, GNN-Ret improves EM by 4.4 and accuracy by $4\%$ on the IIRC dataset. Additionally, GNN-Ret maintains its superiority on the Quality dataset, which requires a comprehensive understanding of the context of an entire story. Notably, GNN-Ret effectively enhances accuracy by $4.8\%$ with $K=10$ on the Quality dataset. This improvement can be attributed to the ability of GNN-Ret to enhance the retrieval prioritization of supporting passages through the incorporation of structural information, thereby improving the comprehension of LLMs in articles. Surprisingly, our proposed one-hop GNN-Ret even slightly outperforms baselines that perform multi-hop retrieval processes and queries of LLMs in accuracy, which highlights the significant potential of leveraging passage relatedness to enhance the retrieval process for answering multi-hop reasoning questions. Furthermore, the results in Table 1 demonstrate that our proposed RGNN-Ret achieves state-of-the-art performance in terms of F1, EM, and accuracy on the three multi-hop reasoning datasets. RGNN-Ret outperforms the best baseline, KGP, by more than $10\%$ in accuracy on 2WikiMQA dataset. This performance improvement can be attributed to our proposed self-critique technique and RGNN. The self-critique technique enables more accurate judgement on the ending of the answering process and the generation of final answers, and the incorporation of RGNN further enhances the retrieval process for each step of answering, which is essential for generating subquestions and subanswers.
4.3 Additional Analysis
Selection of $K$ and $O$. In order to effectively train the GNN and RGNN models, we conduct ablation studies to select hyperparameters $K$ and $O$, which determine the number of relevant nodes to be sampled and the size of the competitive node set, respectively. To assess the impact of different values of $K$ and $O$, we evaluate the accuracy of GNN-Ret and RGNN-Ret on the MuSiQue and IIRC datasets with varying $K$ and $O$. The average accuracy across these two datasets is represented by the green points in Fig. 4. The results demonstrate that both GNN-Ret and RGNN-Ret consistently achieve stable accuracy performance across different settings of $K$ and $O$. Upon analyzing the results, we observe that GNN-Ret achieves the highest accuracy when $K=5$ and $O=25$, while RGNN-Ret performs best with $K=5$ and $O=10$. GNN-Ret adopts a larger value of $O$ since it considers the whole label set during training, while RGNN-Ret only considers the labels for the current step and subsequent steps. Therefore, we select these settings of $K$ and $O$ to train the GNN and RGNN models for experiments reported in Table 1.
(a) GNN-Ret
(b) RGNN-Ret
Figure 4: Accuracy of GNN-Ret and RGNN-Ret with various $K$ and $O$ in Musique and IIRC datasets. The average accuracy of two datasets are displayed in green points.
Effectiveness of components in GNN-Ret. To assess the individual contributions of each proposed component, we conduct ablation studies and present the results in Table 2. We explore an alternative approach by utilizing the mean semantic distance (GNN-Ret (mean)) instead of the minimum semantic distance as the message in equation (3) for the GNN. This modification results in a lower accuracy compared to the use of the minimum semantic distance as the message, which suggests that employing the minimum semantic distance as the message effectively filters out interfering messages from irrelevant neighbors and preserves the most relevant message for the GNN.
Effectiveness of components in RGNN-Ret. We conduct ablation studies to evaluate the effectiveness of each component in RGNN-Ret, and present results in Table 2. We first remove the RGNN and solely performs dense retrieval for our proposed self-critique technique. The results in Table 2 show that it results in an accuracy reduction of $2.1\%$ and $2.7\%$ on MuSiQue and 2WikiMQA datasets, respectively. Next, we incorporate the GNN for the retrieval process of each subquestion, i.e., Self-critique + GNN. This modification improves accuracy on both MuSiQue and 2WikiMQA datasets compared to the settings with dense retrieval, indicating the effectiveness of the GNN in enhancing the retrieval coverage for each subquestion individually. By combining self-critique and RGNN, our proposed method RGNN-Ret achieves accuracy up to $31.3\%$ and $55.6\%$ on MuSiQue and 2WikiMQA datasets, respectively, which demonstrates the effectiveness of RGNN-Ret in enhancing the retrieval by interconnecting with the graph of passages from previous steps. This interconnection enhances the retrieval of supporting passages across multiple steps, leading to more accurate answers in the multi-hop reasoning process.
Method MuSiQue 2WikiMQA
Dense 23.8 42.3
GNN-Ret (mean) 26.3 44.6
GNN-Ret 27.3 46.9
Self-critique + dense 29.2 52.9
Self-critique + GNN 30.8 55.2
RGNN-Ret 31.3 55.6
RGNN-Ret ($Y$) 27.1 53.3
RGNN-Ret ($Y^{t}$) 30.2 54.6
RGNN-Ret 31.3 55.6
Table 2: Ablation studies of components of GNN and RGNN in accuracy with ChatGPT on MuSiQue and 2WikiMQA datasets.
We also explore different settings for the target node sets during the training of RGNN and present results in Table 2. When using the entire node set of ground-truth supporting passages $\mathcal{S}_{Y}$ as the retrieval labels for each subquestion, denoted as RGNN-Ret ($Y$), the accuracy decreases to $27.1\%$ and $53.3\%$ on MuSiQue and 2WikiMQA datasets, respectively. This is because the supporting passages for initial steps are irrelevant to the subquestions at subsequent steps. Including all of them as labels for each subquestion disrupts the training of RGNN. Instead, we employ the corresponding label for each subquestion, denoted as RGNN-Ret ($Y^{t}$). This approach effectively improves the accuracy to $30.2\%$ and $54.6\%$ on the MuSiQue and 2WikiMQA datasets, respectively. Futhremore, our proposed RGNN-Ret method includes label set of supporting passages not only for the current subquestion but also for subsequent subquestions. This comprehensive approach further improves the accuracy to $31.3\%$ and $55.6\%$ on MuSiQue and 2WikiMQA datasets, respectively, which verifies the effectiveness of RGNN-Ret in enhancing the retrieval process for subsequent subquestions.
Statistics of retrieval accuracy. We explore the EM between the retrieved passages and the ground-truth passages and collect the exact-match number of test samples with varying numbers of supporting passages required in questions on MuSiQue and 2WikiMQA datasets. We do not evaluate the statistics of retrieval accuracy for RGNN-Ret without the specific ground-truth supporting passages for each subquestion. The results in Fig. 5 show that our proposed GNN-Ret achieves higher retrieval accuracy compared with Dense on MuSiQue and 2WikiMQA datasets. For instance, GNN-Ret outperforms Dense by 29 and 43 exact-match test samples for the questions that require 2 supporting passages. This demonstrates that the GNN indeed improves retrieval coverage of supporting passages and the retrieval prioritization of supporting passages for subsequent reasoning. We display some case studies to demonstrate advantages of GNN-Ret and RGNN-Ret in Appendix C.
(a) MuSiQue
(b) 2WikiMQA
Figure 5: Exact-match number of test samples with varying numbers of supporting passages required for QA on MuSiQue and 2WikiMQA datasets.
5 Conclusion
In this paper, we proposed GNN-Ret, an effective method to enhance retrieval for QA of LLMs by exploiting the inherent relatedness between passages on a graph of passages. By extending GNN-Ret to multi-hop reasoning questions, we proposed RGNN-Ret, which enhances the retrieval for subquestions through the interconnection between graphs of passages across steps. The experiments clearly demonstrated the superiority of both GNN-Ret and RGNN-Ret over baselines, highlighting the effectiveness of leveraging the relatedness between passages to enhance the retrieval process. From these advantages, we believe that the incorporation of graph-based representations and the exploitation of passage relatedness can open up new avenues of research in the field of LLMs in understanding structural documents and answering complex questions.
Appendix A Training Details
In this section, we supplement the training details of GNN and RGNN as below.
A.1 Training of GNN
Hinge objective for GNN. Recall that the objective of GNN is to reduce the semantic distances of supporting passages more than those of non-target nodes. With the ground-truth set of supporting passages $\mathcal{S}_{Y}$, we first define the average integrated semantic distances of supporting passages $\bar{d}_{Y}^{L}$ and non-target nodes $\bar{d}_{O/Y}^{L}$ as below:
$\bar{d}_{Y}^{L}=\frac{1}{|\mathcal{S}_{Y}|}\sum_{j\in\mathcal{S}_{Y}}h_{j}^{L},\quad\bar{d}_{O/Y}^{L}=\frac{1}{|\mathcal{S}_{O/Y}|}\sum_{o\in\mathcal{S}_{O/Y}}h_{o}^{L},$ (8)
where $S_{O}$ is the competitive node set with the top $O$ smallest semantic distances and $S_{O/Y}$ is the non-target node set by removing the target node set $S_{Y}$ from $S_{O}$, respectively. We only consider the competitive node set since the average integrated semantic distance of all the nodes is significantly large and loses effectiveness in training. With the output average semantic distances of supporting passages and non-target nodes, we formulate the hinge objective function with threshold $r$ as follows:
$\ell=\max(0,r+\bar{d}_{Y}^{L}-\bar{d}_{O/Y}^{L}).$ (9)
We update the parameters of GNN using gradient descent, and the gradient of $\ell$ with respect to $\alpha^{l}$ is given by:
$\frac{\partial\ell}{\partial\alpha^{l}}=\left\{\begin{aligned} 0,\quad\ell<=0\\ f_{\alpha^{l}}(L,\mathcal{S}_{Y}),\quad\ell>0\end{aligned}\right.,$ (10)
where
$f_{\alpha^{l}}(L,\mathcal{S}_{Y})$ $=\frac{\sum_{j\in\mathcal{S}_{Y}}(h_{j}^{l-1}-m_{j}^{l})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{j}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{Y}|}$ (11)
$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}_{O/Y}^{l}}(h_{o}^{l-1}-m_{o}^{l})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{o}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{O/Y}^{l}|}.$
A.2 Training of RGNN
Hinge objective for RGNN. For each subquestion $q^{t}$ at step $t$, we input its semantic distances into GNN and compute hidden state $h_{j}^{l,t}$ and message $m_{j}^{l,t}$ of the $j-$th node for $j\in\{1,\cdots,|\mathcal{V}|\}$ at layer $l$. Recall that we adopt a hinge objective that encourages the average integrated semantic distance of supporting passages that should appear at the current and subsequent steps $t+=\{t,\cdots,T\}$ to be lower than that of the non-target node set for subquestion $q^{t}$ at step $t$. We first define the average semantic distances of supporting passages ${\bar{d}_{Y^{t+}}^{L,t}}$ and non-target nodes $\bar{d}_{O/Y^{t+}}^{L,t}$ at step $t$ as below:
$\bar{d}_{Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{Y^{t+}}|}\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\hat{h}_{j}^{L,t},\quad\bar{d}_{O/Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{O/Y^{t+}}|}\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\hat{h}_{o}^{L,t},$ (12)
where $\mathcal{S}_{O}^{t}$ is the competitive node set with top $O$ smallest semantic distances at step $t$ and $\mathcal{S}_{O/Y^{t+}}^{t}$ is the non-target node set by removing the target node set $\mathcal{S}_{Y^{t+}}^{t}$ from $\mathcal{S}_{O}^{t}$, respectively. We formulate the hinge objective function with threshold $r$ for the RGNN as follows:
$\ell$ $=\frac{1}{T}\sum_{t\in[T]}\ell^{t}=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (13)
The gradient of loss $\ell^{t}$ at step $t$ with respect to $\alpha$ is computed by:
$\frac{\partial\ell^{t}}{\partial\alpha^{l}}=\left\{\begin{aligned} 0,\quad\ell<=0\\ \sum_{\tau=0}^{t}\beta(1-\beta)^{t-\tau}f_{\alpha^{l}}(L,\mathcal{S}^{t}_{Y^{t+}}),\quad\ell>0\end{aligned}\right.,$ (14)
where
$f_{\alpha^{l}}(L,\mathcal{S}^{t}_{Y^{t+}})$ $=\frac{\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}(h_{j}^{l-1,t}-m_{j}^{l,t})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{j}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}^{t}_{Y^{t+}}|}$ (15)
$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}_{O/Y}^{l}}(h_{o}^{l-1,t}-m_{o}^{l,t})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{o}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{O/Y}^{l}|}.$
The gradient of loss $\ell^{t}$ at step $t$ with respect to $\beta$ is computed by:
$\frac{\partial\ell^{t}}{\partial\beta}=\left\{\begin{aligned} 0,\quad\ell<=0\\ f_{\beta}(t,L,\mathcal{S}^{t}_{Y^{t+}}),\quad\ell>0\end{aligned}\right.,$ (16)
where
$f_{\beta}(t,L,\mathcal{S}^{t}_{Y^{t+}})$ $=\frac{\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\sum_{\tau=1}^{t}(h_{j}^{L,\tau}-\hat{h}_{j}^{L,\tau-1})(1-\beta)^{t-\tau}}{|\mathcal{S}^{t}_{Y^{t+}}|}$ (17)
$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\sum_{\tau=1}^{t}(h_{o}^{L,\tau}-\hat{h}_{o}^{L,\tau-1})(1-\beta)^{t-\tau}}{|\mathcal{S}_{O/Y^{t+}}^{t}|}.$
Appendix B Additional Experimental Details
Graph of passages construction. We constructe a large graph of passages to serve as the retrieval corpus for each evaluation dataset. For the Quality dataset, we randomly sample 30 articles with a diverse range of topics as the retrieval corpus. For multi-hop reasoning datasets, we first collect the Wikipedia documents using the provided titles required for answering these 500 questions. For the multi-hop reasoning datasets, we collect Wikipedia documents based on the provided titles required to answer the 500 questions. We chunk the documents into smaller passages and record their sequential order. Passages that are physically adjacent to each other are regarded as structure-related and connected together in the graph. Besides, to find out the keyword-related passages, we also extract the keywords for each passage and then connect those that share the same keywords. Specifically, we prompt ChatGPT (gpt-3.5-turbo-2023-06-01-preview) to extract the Wikipedia keywords from the passages and generate their corresponding links. The links are used to ensure the consistency of the Wikipedia document that use different keywords in passages. The prompt for keyword extraction is given as follows:
Prompt for extracting keywords Instruction: Extract the entities exist in this text and then provide the wikipedia links for the entities exist in this text. Output the entities and their wikipedia links in the list format, e.g., [[’entity1’, ’link1’], [’entity2’, ’link2’]].
<Passage>
The passages share the same links of keywords are considered as keyword-related and connected together in the graph.
Implemental details of training. We initialize $\alpha^{1}=1.0$ and $\beta=1.0$, and use gradient descent to update the GNN and RGNN. For each iteration, we compute the average gradient of all the training samples and update the parameters with a learning rate of 1.0. The training will stop when the absolute value of gradient is smaller than 0.001 or the loss keeps increasing for five consecutive iterations. As there is often the whole label set of supporting passages for the question but not the individual index for each subquestion, we select the one with the lowest semantic distance as the label for each subquestion and then remove this index label from the label set for subsequent subquestions. Since training the RGNN requires ground-truth subquestions, we manually generate subquestions for 5 questions sampled from the preset training data. The training samples of MuSiQue, IIRC, and 2WikiMQA datasets for the RGNN are displayed in Appendix G. We also display the instructions and prompt templates of all the methods in Appendix F.
Appendix C Qualitative Results
We analyze the qualitative results in experiments and demonstrate the effectiveness of GNN-Ret and RGNN-Ret in improving accuracy for QA in this section.
GNN-Ret improves the retrieval coverage of supporting passages. We compare the retrieval process between Dense and GNN-Ret and display the results on Table 3. GNN-Ret can retrieve all of the supporting passages while Dense fails in both of these two cases. For the first question, Dense can retrieve the knowledge that ‘the director of file Hotel By The Hour is Rolf Olsen.’. Since it considers passages in isolation during retrieval, the supporting passage for the inquiry information (‘which country is Rolf Olsen from’) cannot be retrieved with a poor semantic distance. In contrast, GNN-Ret is able to retrieve both of the supporting passages about ‘the director of file Hotel By The Hour is Rolf Olsen’ and ‘Rolf Olsen was an Austrian actor’. Consequently, it can output the correct final answer. This is attributed to the fact that GNN-Ret takes relatedness between passages into account and thus allows the supporting passages for inquiry information to accept the effect of semantic distances from those for background information, thereby improving the retrieval coverage of supporting passages.
Question: Which country the director of film Hotel By The Hour is from? Ground-truth answer: Austria Dense Retrieved passages: (1) Hotel by the Hour (German title:) is a 1970 West German crime film directed by Rolf Olsen and starring Curd Jürgens, Andrea Rau and Corny Collins. (2) … Final answer: Rolf Olsen is from Germany. GNN-Ret Retrieved passages: (1) Hotel by the Hour (German title:) is a 1970 West German crime film directed by Rolf Olsen and starring Curd Jürgens, Andrea Rau and Corny Collins. (2) Rolf Olsen (26 December 1919 – 3 April 1998) was an Austrian actor. Final answer: Rolf Olsen is from Austria. Question: What is the date of death of the performer of song Goodbye Pork Pie Hat? Ground-truth answer: January 5, 1979 Dense Retrieved passages: (1) ‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album. (2) … Final Answer: February 17, 2017 GNN-Ret Retrieved passages: (1)‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album. (2) Charles Mingus Jr. (April 22, 1922 – January 5, 1979) was an American jazz double bassist, pianist, composer and bandleader. Final answer: January 5, 1979
Table 3: Two examples demonstrating effectiveness of GNN-Ret. We highlight the shared keywords, correct answers, and incorrect answers in violet, green, and red.
Question: Which award the performer of song One More Time (Joe Jackson Song) earned? Ground-truth answer: Grammy SelfAsk Generation: Here is the question: Question: Which award the performer of song One More Time (Joe Jackson Song) earned? Are follow up questions needed here: Yes. Follow up: Who is the performer of the song One More Time (Joe Jackson Song)? The performer of the song ""One More Time"" (Joe Jackson Song) is Joe Jackson. So the final answer to the Question is: Joe Jackson RGNN-Ret Generation: Question: Which award the performer of song One More Time (Joe Jackson Song) earned? Next step: What is the name of the performer of the song ""One More Time (Joe Jackson Song)""? Intermediate Answer: The performer of the song ""One More Time"" is Joe Jackson. Next step: What award did Joe Jackson earn for the song ""One More Time""? Intermediate Answer: Joe Jackson earned a Grammy for Best Pop Instrumental Album in 2001 for his album Symphony No. 1. Final answer: Grammy Question: What other notable work did the creator of Shrek make? Ground-truth answer: Doctor De Soto SelfAsk Generation: Question: What other notable work did the creator of Shrek make? Are follow up questions needed here: Yes. Follow up: Who is the creator of Shrek? The creator of Shrek is William Steig. Are follow up questions needed here: Yes. Follow up: What other notable work did William Steig make? William Steig also wrote "Sylvester and the Magic Pebble." So the final answer is: Sylvester and the Magic Pebble RGNN-Ret Generation: Question: What other notable work did the creator of Shrek make? Next step: Who is the creator of Shrek? Intermediate Answer: The creator of Shrek is American author William Steig. Next step: What other notable works did William Steig create besides Shrek? Intermediate Answer: William Steig, the creator of Shrek, also wrote and illustrated the children’s picture book "Doctor De Soto" which was first published in 1982… Final answer: Doctor De Soto
Table 4: Two generation examples of comparison between RGNN-Ret and SelfAsk. We highlight the correct answers and incorrect answers in green and red. We omit the retrieved passages in the table.
RGNN-Ret better determines the terminal of answer process compared with SelfAsk. We analyze the qualitative results for RGNN-Ret and SelfAsk since they have the similar answering procedures by generating subquestions and use them for retrieval. The qualitative results of them are shown in Table 4. For the first question ‘Which award the performer of song One More Time (Joe Jackson Song) earned?’, SelfAsk generates the first-step subquestion and also answer it correctly. It obtains the knowledge that ‘the performer of the song One More Time (Joe Jackson Song) is Joe Jackson’. However, it terminates the answering process at this step and thus outputs the incorrect final answer. In contrast, RGNN-Ret understands that the generated intermediate answers are not sufficient to output the final answer and thus continues the next-step answering. Consequently, it can output the final answer ‘Grammy’. This qualitative comparison between RGNN-Ret and SelfAsk demonstrates the effectiveness of our proposed self-critique technique in determining the termination of the answering process and improving the accuracy for QA.
RGNN-Ret enhances the retrieval process for subquestions. For the second question ‘What other notable work did the creator of Shrek make?’ in Table 4, both SelfAsk and RGNN-Ret are able to correctly answer the first subquestion and generate the second-step subquestion ‘What other notable works did William Steig create?’. However, SelfAsk cannot retrieve the knowledge about ‘other work of William Steig’, which locates in another passage about the book ‘Doctor De Soto’. Using dense retrieval fails to retrieve this passage for QA. In contrast, RGNN-Ret retrieve this passage since it enhances the semantic distance by integrating with the small semantic distances from the previous step. Consequently, RGNN-Ret outputs the correct answer. This qualitative example demonstrates that the RGNN can enhance the retrieval over steps for subquestions and thus improve the accuracy for QA.
Appendix D Limitations and Discussions
Undirected Graph. In this study, our focus lies solely on improving the retrieval process through an undirected graph of passages, with no additional information associated with the edges. While we employ a sampling technique to select relevant nodes and utilize the minimum semantic distance as the message in the GNN, we acknowledge that there is still a possibility of irrelevant passages receiving messages, potentially impacting the retrieval performance. Consequently, the task of selecting an appropriate path on the graph of passages that aligns with the given question remains an area that is yet to be thoroughly explored. In addition, we only consider the structural-related passages that are in the same section. However, some questions may require more complex structural information. For example, when the question is about comparison of two entities in different sections, more complex structural information is needed to connect these corresponding passages to enhance retrieval. Therefore, a more delicate graph of passages for more complex tasks warrant further investigation in future studies.
Costs of graph construction. While our proposed GNN-Ret demonstrates impressive accuracy in QA tasks with a single query of LLMs, it still relies on multiple LLM queries to extract keywords from passages and construct the graph of passages. Fortunately, there are alternative fine-tuned language models specific for keyword extraction [^3] https://huggingface.co/ml6team/keyphrase-extraction-distilbert-inspec [^4] https://github.com/MaartenGr/KeyBERT. These models are more compact and significantly accelerate the graph construction process. Moreover, the flexibility of the graph of passages allows for dynamic modifications, such as adding new passage nodes or removing outdated ones. By maintaining a domain-specific graph of passages, we can effectively address various questions for this domain without the need for reconstructing the graph.
Appendix E Ethics Statement
We hold a firm conviction that our proposed GNN-Ret and RGNN-Ret have the potential to yield substantial positive outcomes in real-world scenarios associated with retrieval-augmented generation (RAG). GNN-Ret and RGNN-Ret significantly enhanced the retrieval coverage of supporting passages, thereby significantly improving the explainability and reliability of LLMs. Given the widespread application of LLMs, we anticipate that the implementation of our proposed retrieval techniques will extend their benefits to an increasing number of individuals, thereby catering to a broader range of needs.
Contrasting the numerous potential benefits of utilizing LLMs, it is crucial to acknowledge the existence of certain negative impacts, such as the generation of hallucinatory outputs and the associated costs. Addressing the issue of hallucinations in LLMs requires ongoing efforts to improve retrieval accuracy and provide comprehensive supporting passages to LLMs. By ensuring that LLMs have access to a wider range of relevant information, the occurrence of hallucinatory responses can be mitigated. Additionally, cost reduction strategies can be implemented by substituting specific steps of GNN-Ret and RGNN-Ret, e.g., keyword extraction. The pretrained small LMs could be an effective way to reduce the costs for applications.
Appendix F Prompts for Experiments
We display prompt templates of all the methods in this section.
F.1 Prompt template for Direct.
The prompt template for Direct is shown as follows:
Prompt template for Direct Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
<Question>
where <Question> indicates the question.
F.2 Prompt templates for Dense, IRCoT, and KGP
Dense, IRCoT, and KGP employ different retrieval methods but the same prompt templates for QA. The prompt template is shown as follows:
Prompt template for Dense / IRCoT / KGP Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
Context: <Context>
Question: <Question>
where <Context> indicates the retrieved passages.
We repeat the question before the length retrieved passages when answering the questions on IIRC and 2WikiMQA datasets for better performance.
F.3 Prompt template for SelfAsk
SelfAsk prompts LLMs to generate the follow-up question and answers it with retrieved passages until generating the final answer. The retrieved passages are included into the prompt when the LLMs are prompted to answer the follow-up question.
Prompt template for SelfAsk Instruction: Your goal is to answer the question step by step following procedures of examples. If no follow up questions are necessary, answer the question in five words directly in format: So the final answer is: xxx.
Here are the examples of how to answer the questions:
<Examples>
Context (Optional): <Context>
Question:
<Question>
We repeat the question before the length retrieved passages when answering the follow-up questions on IIRC and 2WikiMQA datasets for better performance.
F.4 Prompt template for ITER-RETGEN
ITER-RETGEN iteratively answers questions with retrieved passages and uses the generated answer for the next-step retrieval, which continues until the generation of final answer. The prompt template is shown as follows:
Prompt template for ITER-RETGEN Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
Here are the examples of how to answer the questions:
<Examples>
Context: <Context>
Question:
<Question>
Let’s think step by step.
F.5 Prompt template for RGNN-Ret
RGNN-Ret iteratively generates the next-step subquestion, answers the subquestion, and performs self-critique until the generation of final answer. The prompt template of these procedures are shown as follows:
Prompt template for generating next-step subquestions Instruction: Your goal is to ask the next step question logically.
Here are the examples of how to answer the questions:
<Examples>
Question:
<Question>
Prompt template for generating subanswers Instruction: Your goal is to answer the next step question. I will provide you some wikipedia snippets, and you need to answer the next step question by considering the wikipedia snippets.
Context:
<Context>
Question:
<Question>
We repeat the question before the length retrieved passages when answering the next-step subanswer on IIRC and 2WikiMQA datasets for better performance.
Prompt template of self-critique Instruction: You are a wikipedia QA expert. Your goal is to critique whether the intermediate answers are enough to generate the final answer to the question.First analyze if the intermediate answers is enough to generate the final answer step by step logically. Then, if it is enough, output ’Critique: yes’. If not, output ’Critique: no’.
Question:
<Question>
"Analyze if the intermediate answers is enough to generate the final answer to the question step by step logically. Then, if it is enough, output ’Critique: yes’. If not, output ’Critique: no’.
The demonstrations of RGNN-Ret for asking the next-step subquestion on MuSiQue, IIRC, and 2WikiMQA datasets are shown as follows:
Demonstrations of RGNN-Ret for asking next-step subquestions on MuSiQue dataset Question: Who lived longer, Muhammad Ali or Alan Turing?
Next step: How old was Muhammad Ali when he died?
Intermediate answer: Muhammad Ali was 74 years old when he died.
Next step: How old was Alan Turing when he died?
Question: When was the founder of craigslist born?
Next step: Who was the founder of craigslist?
Intermediate answer: Craigslist was founded by Craig Newmark.
Next step: When was Craig Newmark born?
Question: Who was the maternal grandfather of George Washington?
Next step: Who was the mother of George Washington?
Intermediate answer: The mother of George Washington was Mary Ball Washington.
Next step: Who was the father of Mary Ball Washington?
Question: Are both the directors of Jaws and Casino Royale from the same country?
Next step: Who is the director of Jaws?
Intermediate Answer: The director of Jaws is Steven Spielberg.
Next step: Where is Steven Spielberg from?
Intermediate Answer: The United States.
Next step: Who is the director of Casino Royale?
Intermediate Answer: The director of Casino Royale is Martin Campbell.
Next step: Where is Martin Campbell from?
Demonstrations of RGNN-Ret for asking next-step subquestions on IIRC dataset Question: Who lived longer, Muhammad Ali or Alan Turing?
Next step: How old was Muhammad Ali when he died?
Intermediate answer: Muhammad Ali was 74 years old when he died.
Next step: How old was Alan Turing when he died?
Question: When was the founder of craigslist born?
Next step: Who was the founder of craigslist?
Intermediate answer: Craigslist was founded by Craig Newmark.
Next step: When was Craig Newmark born?
Question: Was the city where Eva was born the capital of its country?
Next step: Where was Eva born?
Intermediate answer: Eva was born in Berlin, Germany.
Next step: Is Berlin the capital of Germany?
Question: Was Ryuji Yamakawa a good solo wrestler?
Next step: Who is Ryuji Yamakawa?
Intermediate answer: Ryuji Yamakawa is a Japanese professional wrestler.
Next step: Was Ryuji Yamakawa a good solo wrestler?
Demonstrations of RGNN-Ret for asking next-step subquestions on 2WikiMQA dataset Question: Who lived longer, Muhammad Ali or Alan Turing?
Next step: How long did Muhammad Ali live?
Intermediate answer: Muhammad Ali was 74 years old when he died.
Next step: How long did Alan Turing live?
Question: Who was the paternal grandfather of Princess Alexandrine Of Prussia (1842-1906)?
Next step: Who was the father of Princess Alexandrine Of Prussia (1842-1906)?
Intermediate answer: Princess Alexandrine Of Prussia (1842-1906) was the daughter of Prince Albert of Prussia.
Next step: Who was the father of Prince Albert of Prussia?
Question: Who is the mother of the composer of film 404 (Film)?
Next step: Who is the composer of film 404 (Film)?
Intermediate answer: The composer of film 404 (Film) is Ilayaraja.
Next step: Who is the mother of Ilayaraja?
Question: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality?
Next step: What is the nationality of the director of Across The Badlands?
Intermediate answer: The director of Across The Badlands is American.
Next step: What is the nationality of the director of A Gutter Magdalene?
Intermediate answer: The director of A Gutter Magdalene is American.
Next step: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality?
Question: What is the home stadium of the team that Asprey hit a hat trick against on 16 January 1961?
Next step: Which team did Asprey hit a hat trick against on 16 January 1961?
Intermediate answer: Asprey hit a hat trick against Charlton Athletic on 16 January 1961.
Next step: What is the home stadium of Charlton Athletic?
Appendix G Training data for RGNN
We manually generate subquestions for 5 questions sampled from the preset training data for MuSiQue, IIRC, and 2WikiMQA datasets. The training samples are shown as follows:
Training data of MuSiQue for RGNN 1. Question: Why did Roncali leave the place of death of the creator of Malchiostro Annunciation?
Subquestions: Who is the creator of Malchiostro Annunciation? Where did Titian die? Why did Roncali leave Venice?
2. Question: Where did who argued that the country of citizenship of Victor Denisov had itself beome an imperialist power declare that he would intervene in the Korean conflict?
Subquestions: What is the country of citizenship of Victor Denisov? Who argued that Russia had itself become an imperialist power? Where did Mao Zedong declare that he would intervene in the Korean conflict?
3. Question: What military overran much of Erich Zakowski’s place of birth?
Subquestions: What is the place of birth of Erich Zakowski? What military overran East Prussia?
4. Question: How many people were in British Colonies where does the london broil cut come from ?
Subquestions: Where does the london broil cut come from? How many people were in North American?
5. Question: When was the country established that lies immediately north of the Persian Gulf and the region where the country containing Urim is located?
Subquestions: What is the region containing Urim? Where is the region that Iraq is located? What is the country that lies immediately north of the Persian Gulf and the Middle East? When was Iran established?
Training data of IIRC for RGNN 1. Question: Was the city where Eva was born the capital of its country?
Subquestions: Where was Eva born? Is Berlin the capital of its country?
2. Question: In what state did Galambos attend high school?
Subquestions: What high school did Galambos attend? In what state is Athens High School located?
3. Question: What was the previous name of the team that Feng started playing with in 1999?
Subquestions: What team did Feng start playing with in 1999? What was the previous name of Chongqing Longxin?
4. Question: How many years after the first Marvel Cinematic Universe film came out was Black Panther released?
Subquestions: When was the first Marvel Cinematic Universe film released? When was Black Panther released? How many years after Iron Man came out was Black Panther released?
5. Question: Who was the first draft pick the year Damarius Bilbo went undrafted?
Subquestions: What year did Damarius Bilbo go undrafted? Who was the first draft pick in 2006?
Training data of 2WikiMQA for RGNN 1. Question: Which film came out earlier, Subliminal Seduction or Australia Marches With Britain?
Subquestions: When did Subliminal Seduction come out? When did Australia Marches With Britain come out?
2. Question: Who is the father-in-law of Deuteria?
Subquestions: Who is the husband of Deuteria? Who is the father of Eusebio?
3. Question: Who is the mother of the composer of film 404 (Film)?
Subquestions: Who is the composer of film 404 (Film)? Who is the mother of Ilayaraja?
4. Question: Which country the composer of film Sergeant Hassan is from?
Subquestions: Who is the composer of film Sergeant Hassan? Which country is Tamer Karaoğlu from?
5. Question: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality?
Subquestions: What is the director of Across The Badlands? Where is Fred F. Sears from? What is the director of A Gutter Magdalene? Where is George Melford from?

### v2

@@
-Graph Neural Network Enhanced Retrieval for Question Answering of LLMs
+Graph Neural Network Enhanced Retrieval for Question Answering of Large Language Models
 Abstract
 […]
+Graph Neural Network Enhanced Retrieval for Question Answering of Large Language Models
+Zijian Li12thanks: Work done during an internship at Microsoft Research Asia.
+, Qingyan Guo23, Jiawei Shao1, Lei Song2, Jiang Bian2, Jun Zhang1thanks: Corresponding Author.
+, Rui Wang2† 1Hong Kong University of Science and Technology 2Microsoft Research 3Tsinghua University {zijian.li,jiawei.shao}@connect.ust.hk, gqy22@mails.tsinghua.edu.cn, {lesong,jiabia,ruiwa}@microsoft.com, eejzhang@ust.hk
 1 Introduction
-Large language models (LLMs) continue to struggle with factual errors when encountering knowledge intensive questions [7, 19, 9].
-Although retrieval-augmented LLMs [16, 15] have improved factuality and precision of question answering by including relevant passages, there remains a persistent challenge in accurately capturing all the relevant supporting passages when encountering complex knowledge-intensive questions.
+Large language models (LLMs) continue to struggle with factual errors when encountering knowledge intensive questions (Huang et al., 2023; Mallen et al., 2022; Ji et al., 2023).
+Although retrieval-augmented LLMs (Lewis et al., 2020) have improved factuality and precision of question answering by including relevant passages, there remains a persistent challenge in accurately capturing all the supporting passages when encountering complex knowledge-intensive questions.
 This limitation can be attributed to the inherent information asymmetry in complex questions.
-In particular, the questions tend to consist of elaborate background details, leaving only a small portion dedicated to specific inquiries.
-As an example, in question: ‘Why did crime rise on Mars after the Mafia’s arrival?’, the majority part delves into the background (i.e., ‘crime rise …… Mafia’s arrival’) with only a few words requesting the reason (i.e., ‘Why’), which consequently retrieves passages that focus on the details of crime rising instead of the reason for it.
+In particular, the questions tend to consist of elaborated background details, leaving only a small portion dedicated to specific inquiries.
+As an example, in question: ‘Why did crime rise on Mars after the Mafia’s arrival?’, the majority part delves into the background (i.e., ‘crime rise …… Mafia’s arrival’) with only a few words requesting the reason (i.e., ‘Why’), which consequently retrieves passages on details of crime rising instead of the reason for it.
 This phenomenon also frequently arises in multi-hop reasoning questions.
@@
 1, it becomes apparent that while we can retrieve the knowledge that the performer of song Left & Right is D’Angelo, his birthplace remains absent.
-Previous works [30, 23, 35] have attempted to address this issue by incorporating multi-hop reasoning or question rewriting into retrieval processes, which enables them to retrieve information based on prior reasoning outcomes.
-LLMs, however, often generate plausible reasons and incorrect rewriting questions without accurate prior domain knowledge of the given question [7, 36], thus affecting the subsequent retrieval process.
+Previous works Trivedi et al.
+(2023); Press et al.
+(2022); Yao et al.
+(2022) have attempted to address this issue by incorporating multi-hop reasoning or question rewriting into retrieval processes, which enables them to retrieve information based on prior reasoning outcomes.
+LLMs, however, often generate plausible reasons and incorrect rewriting questions without accurate prior domain knowledge of the given question Huang et al.
+(2023); Zhang et al.
+(2023), thus affecting the subsequent retrieval process.
 Figure 1: Overview of comparison between dense retrieval and GNN-Ret.
@@
 By considering the relatedness between passages, GNN-Ret can retrieve all the supporting passages for QA.
-One reason why existing methods struggle to handle the information asymmetry is their tendency to consider passages in isolation [14] and retrieve the supporting passages based on semantic distances, making it difficult to retrieve all the supporting passages, especially those containing only few words for inquiry.
+One reason why existing methods struggle to handle the information asymmetry is their tendency to consider passages in isolation Karpukhin et al.
+(2020a) and retrieve the supporting passages based mainly on semantic distances, making it difficult to retrieve all the supporting passages, especially those containing only few words for inquiry.
 However, the supporting passages of background and inquiry information are usually correlated.
@@
 The key challenge lies in how to effectively leverage the passage of graphs to enhance retrieval coverage.
-Graph neural networks (GNNs) are neural networks tailored to analyze graph data and adeptly grasp the relationships between nodes and edges [26].
+Graph neural networks (GNNs) are neural networks tailored to analyze graph data and adeptly grasp the relationships between nodes and edges Scarselli et al.
+(2008).
 Thus, we propose to leverage a GNN to enhance the retrieval process by effectively capturing the relatedness between passages and name this method GNN-Ret.
 The GNN facilitates the integration of semantic distances between related passages, thus enabling the semantic distances of passages containing background information to impact the retrieval of passages relevant to the inquiry.
-Figure 2: Accuracy and average query times of LLMs for our proposed methods and baselines on 2WikiMQA dataset.
 To address multi-hop reasoning questions, we further propose a retrieval method, named RGNN-Ret, which leverages a recurrent graph neural network (RGNN) to enhance the retrieval process at each step by integrating the retrievals from previous steps through the interconnected graphs of passages.
 This boosts the retrieval coverage for supporting passages over steps even when LLMs generate incorrect reasons or subquestions for retrieval.
-Through the experiments on four benchmark datasets, we demonstrate the effectiveness of GNN in enhancing retrieval coverage for supporting facts and thus improving accuracy for QA tasks.
+Figure 2: Accuracy and average number of LLM queries for our proposed methods and baselines on 2WikiMQA.
+Through the experiments on four benchmark datasets, we demonstrate the effectiveness of GNNs in enhancing retrieval of supporting passages and thus improving accuracy for QA.
 For example, as shown in Fig.
-2, our proposed GNN-Ret with a single query of LLMs significantly outperforms baselines in accuracy on the 2WikiMQA dataset [6], including those methods that require multiple queries of LLMs, which demonstrates the effectiveness of GNNs in improving retrieval coverage of supporting passages.
+2, our proposed GNN-Ret with a single query of LLMs significantly outperforms baselines in accuracy on 2WikiMQA Ho et al.
+(2020), including those methods that require multiple queries of LLMs.
 Moreover, by extending GNN to multi-hop reasoning questions, our proposed RGNN-Ret achieves state-of-the-art accuracy.
@@
 Retrieval-augmented LLM.
-QA [31] is a task that often requires external and up-to-date knowledge sources to answer factoid-based questions.
-Retrieve-and-read is the basic framework for these questions [5, 37].
-For retrieval, sparse retrievers, e.g., TF-IDF and BM25 [24], or dense retrievers, e.g., DPR [13] and Contriever [8], are used to compute lexical distances or semantic distances between passages and the question.
-The passages with smaller distances are retrieved, which are then prefixed with the question for factual answering [16].
+QA Voorhees et al.
+(1999) is a task that often requires external and up-to-date knowledge sources to answer factoid-based questions.
+Retrieve-and-read is the basic framework for these questions Gao et al.
+(2023); Zhu et al.
+(2023).
+For retrieval, sparse retrievers, e.g., TF-IDF and BM25 Robertson et al.
+(2009), or dense retrievers, e.g., DPR Karpukhin et al.
+(2020b) and Contriever Izacard et al.
+(2021), are applied to compute lexical distances or semantic distances between passages and the question.
+The passages with smaller distances are retrieved, which are then prefixed with the question for factual answering Lewis et al.
+(2020).
 However, this framework treats passages as isolated units, making it difficult to retrieve all the supporting passages for the question in a single step.
-To address this limitation, many approaches have been proposed to integrate the retrieval and reasoning processes to improve the retrieval of supporting passages [5, 27, 3, 30, 11, 35, 23].
-They first prompt LLMs to generate the next-step reason or subquestion and then use it to guide retrieval or knowledge search for QA.
-However, a challenge in these methodologies pertains to the potential generation of hallucinated reasons or erroneous subquestions by LLMs without accurate domain knowledge of this question, leading to degraded subsequent retrievals [7].
+To address this limitation, many approaches have been proposed to integrate the retrieval and reasoning processes to improve the retrieval of supporting passages Gao et al.
+(2023); Shao et al.
+(2023); Feng et al.
+(2024); Trivedi et al.
+(2023); Jiang et al.
+(2023b); Yao et al.
+(2022); Press et al.
+(2022).
+They first prompt LLMs to generate the next-step reason or subquestion and then use it to guide retrieval for QA.
+However, a challenge in these methodologies pertains to the potential generation of hallucinated reasons or erroneous subquestions by LLMs without accurate domain knowledge of this question, leading to degraded subsequent retrievals Huang et al.
+(2023).
 Graph-enhanced LLM.
-As structured, explicit, and editable representations of knowledge, knowledge graphs (KGs) have been effectively used for boosting retrieval coverage [20] and enhancing reasoning capabilities of LLMs [17, 34, 1, 32, 22, 2].
-Instead of naively combining KGs with LLMs, many works directly prompt LLMs to perform reasoning on KGs [28, 10, 18].
-They first identified the entities on the knowledge graph and then search reasoning paths accordingly [28, 18, 10].
+As structured, explicit, and editable representations of knowledge, knowledge graphs (KGs) have been effectively used for boosting retrieval coverage Min et al.
+(2019) and enhancing reasoning capabilities of LLMs Li et al.
+(2023a); Xie et al.
+(2022); Baek et al.
+(2023); Wang et al.
+(2023); Park et al.
+(2023); Ding et al.
+(2024); Jin et al.
+(2023); Li et al.
+(2023b); Xiong et al.
+(2024); He et al.
+(2024).
+Instead of naively combining KGs with LLMs, many works directly prompt LLMs to perform reasoning on KGs Sun et al.
+(2024); Jiang et al.
+(2023a); LUO et al.
+(2024).
+They first identified the entities on the knowledge graph and then search reasoning paths accordingly Sun et al.
+(2024); LUO et al.
+(2024); Jiang et al.
+(2023a).
 However, KG-based methods encounter a limitation that they can only handle questions that are effectively represented as KGs.
 They may struggle with questions that necessitate a comprehensive understanding of lengthy contextual information, such as the mentioned example: the cause of crime increasing.
-Rather than KGs, a recent study [12] attempts to construct a graph of passages and perform reasoning on this graph.
+Rather than KGs, a recent study Jin et al.
+(2024) attempts to construct a graph of passages and perform reasoning on this graph.
 It connects documents according to some specific words and performs reasoning on this graph for QA.
 In contrast, we utilize the graph of passages to enhance the retrieval process.
-Another related work, KGP [33], performs retrieval on the graph and search related passages.
-However, this approach does not consider the integration between related passages and may not effectively enhance retrieval coverage for the supporting passages with poor semantic similarity.
+Another related work, KGP Wang et al.
+(2024), performs retrieval on the graph and search related passages.
+However, this approach does not consider the integration between related passages and may not effectively enhance retrieval for the supporting passages with poor semantic similarity.
 3 Graph Neural Network Enhanced Retrieval for QA
-In this section, we aim to enhance the retrieval coverage of supporting passages by leveraging the relatedness between passages for QA tasks.
-Specifically, we construct a graph of passages by connecting passages that are structural-related and keyword-related, where each node represents a passage (Section 3.2).
+The QA system typically adopts a retrieve-and-read routine Lewis et al.
+(2020): the retriever computes semantic distances between passages and the question in the embedding space, finds out relevant passages from the whole corpus, and then the reader (LLM) generates the answer based on them.
+Considering the information asymmetry of complex questions, directly using the initial semantic distances of passages for retrieval is difficult to retrieve all the supporting passages.
+Therefore, we propose to process by taking the relatedness between passages into account.
+Specifically, we construct a graph of passages (GoPs) by connecting related passages (Section 3.1).
 To exploit the relatedness between passages to enhance retrieval, we utilize a GNN to enable integration between related passages (Section 3.2).
-For multi-hop questions, we further leverage an RGNN to establish relationship between graphs over steps to enhance retrieval for the multi-hop answering process (Section 3.4).
-3.1 Problem Description
-The QA system typically adapts a retrieve-and-read routine [16]: the retriever computes semantic distances $D_{q}$ between passages and question $q$ in the embedding space, finds out the relevant passages $P_{q}$ from the whole corpus $P$, and then the reader (LLM) generates the answer $a$ based on them.
-This process can be formulated as follows:
-$D_{q}=\{\phi(q,p)|p\in P\},\quad P_{q}=\text{Retrieve}(q,D_{q}),\quad a=\text{Read}(q,P_{q}),$ (1)
-where $\phi(q,p)$ measures the semantic distance between question $q$ and passage $p$.
-A smaller value of $\phi(q,p)$ indicates higher similarity between $q$ and $p$.
-Considering the information asymmetry of complex questions, directly using the initial semantic distances of passages $D_{p}$ for retrieval is difficult to retrieve all the supporting passages.
-Therefore, we propose to process $D_{q}$ by taking the relatedness between passages into account and thus enhance the retrieval for QA.
-3.2 Graph of Passages
+For multi-hop questions, we further leverage an RGNN to establish relationship between graphs over steps to enhance retrieval over answering processes (Section 3.3).
+3.1 Graph of Passages (GoPs)
 Establish relationships between related passages.
-Existing retrieval methods perform retrieval merely relying on semantic distances between passages and the question [5, 37], which may fail to retrieve some supporting passages with large semantic distances to the question due to the information asymmetry.
-However, these supporting passages are usually related to those with small semantic distances to the question.
-They can be structural-related, e.g., the happened event (‘crime rising’) and its corresponding reasons are located in the same section.
+Existing retrieval methods rely on semantic distances between passages and the question Gao et al.
+(2023); Zhu et al.
+(2023), which may fail to retrieve some supporting passages with large semantic distances, but related to those with small semantic distances to the question, due to the information asymmetry.
+These passages can be structural-related, e.g., the happened event (‘crime rising’) and its corresponding reasons are located in the same section.
 As shown in Fig.
 1, they can also be keyword-related, e.g., the supporting passages for background information (‘performer of song Left &\&& Right’) and inquiry information (‘the birthplace of D’Angelo’) share the same keyword (‘D’Angelo’).
-Therefore, we establish relationship between passages using structural information and shared keywords to construct the graph of passages.
-Specifically, when chunking the documents, we record the order of passages and connect passages that are physically next to each other in the documents.
-In addition, we extract keywords from passages by prompting LLMs and connect passages containing the same keywords [20, 33].
+Therefore, we propose to establish relationship between passages using structural information and shared keywords to construct the graph of passages.
+Specifically, when chunking the documents, we record the order of passages and connect passages that are physically next to each other in documents.
+In addition, we extract keywords from passages by prompting LLMs and connect passages containing the same keywords Min et al.
+(2019); Wang et al.
+(2024).
 As such, the related passages are connected, having the potential to facilitate retrieval.
-3.3 GNN Enhanced Retrieval on Graph of Passages
-After establishing connections between related passages, we obtain a graph of passages, with each passage as a node.
-When conducting the retrieval process, we first compute the semantic distances between passages and the question $q$ [5, 37].
-The semantic distances of related passages are connected according to the graph of passages.
-To take relatedness between passages into account, we utilize a GNN to process semantic distances based on the graph of passages and obtain the integrated semantic distances, which benefits the retrieval of supporting passages.
+3.2 GNN Enhanced Retrieval on GoPs
+After establishing connections between related passages, we obtain a GoPs, with each passage as a node.
+When conducting the retrieval process, we first compute the semantic distances between passages and the question $q$ Gao et al.
+(2023); Zhu et al.
+(2023).
+The semantic distances of related passages are connected according to the GoPs.
+To take relatedness between passages into account, we utilize a GNN to process semantic distances based on GoPs and obtain the integrated semantic distances, which benefits the retrieval of supporting passages.
 Graph neural network.
-A graph is defined as an ordered pair $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{G}$ is the set of nodes $v_{i}$ and $\mathcal{E}$ is the set of edges $e_{ij}$ connecting two nodes $v_{i}$ and $v_{j}$.
+A graph is defined as an ordered pair $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ is the set of nodes $v_{i}$ and $\mathcal{E}$ is the set of edges $e_{ij}$ connecting two nodes $v_{i}$ and $v_{j}$.
 We define $h_{i}^{l}$ as the hidden state (integrated semantic distance) of the $i-$th node for $i\in\{1,\cdots,|\mathcal{V}|\}$ at layer $l$.
@@
 The set of nodes $\mathcal{N}_{v_{i}}$ stands for the neighbors of node $v_{i}$.
-For an $L-$layer GNN [26], the hidden state of each node $v_{i}\in\mathcal{V}$ is updated through iterative interactions with neighbors.
+For an $L-$layer GNN Scarselli et al.
+(2008), the hidden state of each node $v_{i}\in\mathcal{V}$ is updated through iterative interactions with neighbors.
 Specifically, at layer $l$, the received messages $m_{i}^{l}$ of node $v_{i}$ are calculated using the hidden states of its neighbors.
-Then the received messages $m_{i}^{l}$ and the previous-layer hidden state $h_{i}^{l-1}$ are utilized to compute the hidden state $h_{i}^{l}$ at layer $l$, which are formulated as follows:
-$m_{i}^{l}=f_{m}^{l}(h_{j}^{l-1}|\forall j\in\{1,\cdots,|\mathcal{N}_{v_{i}}|\}),\quad h_{i}^{l}=f_{h}^{l}(m_{i}^{l},h_{i}^{l-1}),$ (2)
-where functions $f_{m}^{l}(\cdot)$ and $f_{h}^{l}$ aggregate neighbors’ information and update the hidden state, respectively.
+Then the received messages $m_{i}^{l}$ and the previous-layer hidden state $h_{i}^{l-1}$ are utilized to compute the hidden state $h_{i}^{l}$ at layer $l$.
 By processing hidden states for all the passages layer by layer, we obtain the integrated semantic distance $h_{i}^{L}$ for each node $v_{i}\in\mathcal{V}$.
-We elaborately introduce the GNN for retrieval as below.
+We elaborately introduce GNNs for retrieval as below.
 Minimum semantic distance as message.
 Since each passage maintains and shares many keywords with other passages, each passage has a large number of neighbors.
-Some of them are not relevant to the question, contributing to distracting or misleading information.
+Some of them are not relevant to the question, contributing to large semantic distances and misleading information.
 Therefore, each node $v_{i}$ only receives the message from the neighbor that has the minimum (integrated) semantic distance to the question at layer $l$.
 The received message of node $v_{i}$ is thus formulated as $m_{i}^{l}=\min_{j\in\mathcal{N}_{v_{i}}}h_{j}^{l-1}$.
+Figure 3: Illustration of GNN-Ret and RGNN-Ret.
 Relevant nodes sampling.
-Given the large scale of the graph constructed by plenty of passages, it is computationally inefficient to propagate information across all nodes.
+Given the large scale of GoPs with plenty of nodes, it is computationally inefficient to propagate information across all of them.
 Besides, allowing irrelevant nodes to pass messages to their neighbors may affect the retrieval process.
@@
 $h_{i}^{l}=\alpha^{l}h_{i}^{l-1}+(1-\alpha^{l})m_{i}^{l}.$ (2)
-The nodes without received message maintain their hidden states of the previous layer.
+The nodes without received messages maintain their hidden states of the previous layer.
 For the GNN with $L$ layers, we compute the hidden state layer by layer and then obtain the integrated semantic distance $h_{i}^{L}$ for each node $v_{i}$.
 With GNN, the supporting passages for inquiry information integrate with the small semantic distance from those for background information and thereby obtain smaller integrated semantic distances, promoting them to be retrieved.
-Figure 3: Illustration of GNN-Ret and RGNN-Ret.
 Hinge objective for GNN.
@@
 The information asymmetry phenomenon also frequently arises in complex multi-hop reasoning questions, as analyzed in the Introduction section.
-Although recent works have incorporated multi-hop reasoning or question rewriting into retrieval [35, 30, 27], LLMs may generate hallucinated reasons or incorrect subquestions in the absence of prior knowledge in this domain [7].
+Although recent works have incorporated multi-hop reasoning or question rewriting into retrieval Yao et al.
+(2022); Trivedi et al.
+(2023); Shao et al.
+(2023), LLMs may generate hallucinated reasons or incorrect subquestions due to the absence of prior knowledge to the question Huang et al.
+(2023).
 These hallucinated reasons or incorrect subquestions fail to retrieve supporting passages when only considering the retrieval process in isolation for each step.
@@
 Otherwise, it will be requested to generate the next-step subquestion.
-Compared with the prior work, SelfAsk [23], we focus on whether existing evidences is satisfactory for answering the question instead of whether the next-step subquestion is needed.
+Compared with SelfAsk Press et al.
+(2022), we omit the subquestions and focus on whether existing evidences are satisfactory for answering the question when determining the termination.
 This self-critique technique decomposes the question over steps and enables a more accurate answer to the question.
@@
 3, we utilize an RGNN to establish relationships between the semantic distances of subquestions.
-Specifically, with the subquestion $q^{t}$ at step $t$, we first compute $h_{i}^{L,t}$ for each node $v_{i}\in\mathcal{V}$ using the aforementioned $L$-layer GNN with Eq.
-(3) and (4).
+Specifically, with the subquestion $q^{t}$ at step $t$, we first compute $h_{i}^{L,t}$ for each node $v_{i}\in\mathcal{V}$ using the aforementioned $L$-layer GNN.
 To consider the effect of semantic distances from previous subquestions, we integrate $h_{i}^{L,t}$ with the previous-step integrated semantic distance to compute the integrated semantic distance $\hat{h}_{i}^{L,t}$ at step $t$ using the parameter $\beta$.
 This can be formulated as follows:
-$\hat{h}_{i}^{L,t}=\beta h_{i}^{L,t}+(1-\beta)\hat{h}_{i}^{L,t-1},\quad(t>0),$ (6)
-where $\hat{h}_{i}^{L,t-1}$ is the integrated semantic distance from the previous step $t-1$ and $\hat{h}_{i}^{L,0}=h_{i}^{L,0}$.
+$\hat{h}_{i}^{L,t}=\beta h_{i}^{L,t}+(1-\beta)\hat{h}_{i}^{L,t-1},\quad(t>1),$ (4)
+where $\hat{h}_{i}^{L,t-1}$ is the integrated semantic distance from the previous step $t-1$ and $\hat{h}_{i}^{L,1}=h_{i}^{L,1}$.
 The integrated semantic distance $\hat{h}_{i}^{L,t}$ will be used to integrate with ${h}_{i}^{L,t+1}$ of node $v_{i}$ at the next step $t+1$ recurrently until the last step of the answering process $T$.
@@
 The hinge objective function with threshold $r$ for the RGNN is formulated as follows:
-$\ell$ $=\frac{1}{T}\sum_{t\in[T]}\ell^{t}=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (7)
+$\ell=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (5)
 By employing this hinge objective, the RGNN takes into account the retrieval of subsequent answering steps and proactively reduces the semantic distances associated with them.
 These small semantic distances are then transferred to the next subquestion and enhance the retrieval of the corresponding supporting passages.
-We update the RGNN (parameters $\alpha^{l}$ and $\beta$) using the gradient descend, and more training details are elaborated in Appendix A.2.
+We update the RGNN (parameters $\alpha^{l}$ and $\beta$) using the gradient descend, which are elaborated in Appendix A.2.
 4 Experiments
-In this section, we aim to conduct experiments to demonstrate the effectiveness of GNN-Ret and RGNN-Ret as follows: 1) Our proposed GNN-Ret and RGNN-Ret achieve superior accuracy on various QA datasets (Table 1).
-2) We conduct ablation studies to select hyperparameters $K$ and $O$ for training GNN and RGNN (Fig.
-4) and verify the effectiveness of each component in GNN-Ret and RGNN-Ret (Table 2).
-3) GNN-Ret improves the retrieval coverage of supporting passages (Fig.
-5).
+In this section, we conduct experiments to demonstrate the effectiveness of GNN-Ret and RGNN-Ret, including 1) superiority of GNN-Ret and RGNN-Ret on various QA datasets (Table 1), 2) ablation studies to verify the effectiveness of components in GoPs, GNN-Ret, and RGNN-Ret (Table 2 and 3), and 3) superiority of GNN-Ret on open-sourced LMs (Table 4).
+Additional analysis on hyperparameter selection, statistics of retrieval accuracy, token number for retrieval, multi-layer GNN-Ret, extension of RGNN-Ret to other multi-hop answering methods, and qualitative results are discussed in Appendix C, D, E, F, G, H, respectively.
 4.1 Experimental Setups
 Evaluation datasets.
-We measure all the methods on four different QA datasets, including multi-hop Wikipedia reasoning datasets: (1) MuSiQue [29], (2) IIRC [4], (3) 2WikiMQA [6], and a single-hop multi-choice QA dataset: (4) Quality [21].
-For each multi-hop dataset, we randomly sample 500 questions from the development set, which provides the ground-truth supporting passages.
-We use 20 of them to train the GNN and select the rest as test data.
-We use all the Wikipedia documents of these 500 questions as the retrieval corpus, which is significantly larger than that provided by datasets [^1] Prior works often use 10–20 paragraphs or documents as the retrieval corpus for each question [27, 33]..
-For each document or article, we split it into multiple passages with a maximum token length of 200 [25].
+We measure all the methods on four different QA datasets, including multi-hop Wikipedia reasoning datasets: (1) MuSiQue Trivedi et al.
+(2022), (2) IIRC Ferguson et al.
+(2020), (3) 2WikiMQA Ho et al.
+(2020), and a single-hop multi-choice QA dataset: (4) Quality Pang et al.
+(2021).
 As evaluation metrics, we calculate the F1 score, exact match (EM) and accuracy (Acc) for multi-hop reasoning datasets.
-We use the ChatGPT (gpt-3.5-turbo-2023-06-01-preview) to evaluate if the prediction matches with the gold answer.
-For Quality, we randomly sample 30 articles with overall corresponding 561 questions and use these articles as the retrieval corpus.
-Since Quality is not a multi-hop dataset, we only validate the accuracy performance of GNN-Ret on it.
+We use the ChatGPT (gpt-3.5-turbo-2023-06-01-preview) to evaluate if the prediction matches with the gold answer, following Shao et al.
+(2023); Wang et al.
+(2024).
+Since Quality is not a multi-hop dataset, we only validate accuracy performance of GNN-Ret on it.
 Baselines.
 We compare GNN-Ret and RGNN-Ret with the following baselines: (1) Direct answers questions without retrieved passages.
-(2) Dense [14] answers questions with retrieved passages using dense retrieval.
-(3) SelfAsk [23] prompts LLMs to generate the follow-up question and answers it with retrieved passages until generating the final answer.
-(4) ITER-RETGEN [27] iteratively answers questions with retrieved passages and uses the generated answer for the next-step retrieval until the generation of final answer.
-(5) IRCoT [30] iteratively prompts LLMs to generate chains of thoughts with retrieved passages and retrieves with the generated reasons until reaching the maximum token number.
+(2)We use retrievers bm25 Robertson et al.
+(2009), DPR Karpukhin et al.
+(2020a), and SentenceBert [^1] https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1 (SBERT) Reimers and Gurevych (2019) to retrieve passages without considering relatedness between passages for QA.
+(3) SelfAsk Press et al.
+(2022) prompts LLMs to generate the follow-up question and answers it with retrieved passages until generating the final answer.
+(4) ITER-RETGEN Shao et al.
+(2023) iteratively answers questions with retrieved passages and uses the generated answer for the next-step retrieval until the generation of final answer.
+(5) IRCoT Trivedi et al.
+(2023) iteratively prompts LLMs to generate chains of thoughts with retrieved passages and retrieves with the generated reasons until reaching the maximum token number.
 All the retrieved passages are then used to generate the final answer.
-(6) KGP [33] first searches seed passages using the question and then prompts LLMs to generate the needed evidence to retrieve the other relevant passages among neighbors of seed nodes using semantic distances.
-Implementation details.
-We use the ChatGPT (gpt-3.5-turbo-2023-06-01-preview) as the LLM backbone for experiments and adopt a temperature of 0 to remove the effect of random sampling.
-We select mpnet-base-cos-v1222https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1 as the embedding model to compute the semantic distance $=1-\text{cosine similarity}$ following [25] for all the methods.
-We set up a maximum token number of 3500 for retrieval to leave the space to instruction and demonstrations, and all the methods retrieve the semantically similar passages until reaching the maximum token number.
-We adopt a one-layer GNN for our proposed methods GNN-Ret and RGNN-Ret.
-We set up $\alpha^{1}=0.5$ for GNN-Ret and $\alpha^{1}=0.5$ and $\beta=0.9$ for RGNN-Ret when there is no training.
-When training the RGNN, we set up $r=0.01$, $K=5$, and $O=25$ for GNN-Ret (w.
-train) and $r=0.01$, $K=5$, and $O=10$ for RGNN-Ret (w.
-train), respectively, according to the ablation studies in Fig.
-4.
-We do not train the parameters for the Quality dataset without the ground-truth labels of supporting passages.
-Since training the RGNN requires ground-truth subquestions, we manually generate subquestions for 5 questions sampled from the preset training data.
-More implementation details are present in Appendix B.
+(6) KGP Wang et al.
+(2024) first searches seed passages using the question and then prompts LLMs to generate the needed evidence to retrieve the other relevant passages among neighbors of seed nodes using semantic distances.
 Methods MuSiQue IIRC 2WikiMQA Quality
@@
 One-hop Retrieval
-Dense 28.6 13.5 23.8 31.6 17.5 40.4 43.4 31.0 42.3 - - 53.4
+bm25 22.2 10.2 18.8 27.3 15.4 31.7 33.8 23.8 34.6 - - 47.0
+DPR 23.8 10.8 17.5 29.9 16.3 35.6 36.2 24.4 36.3 - - 52.5
+SBERT 28.6 13.5 23.8 31.6 17.5 40.4 43.4 31.0 42.3 - - 53.4
 GNN-Ret (K=5) 31.6 16.5 27.1 32.2 17.6 44.0 47.7 32.7 44.8 - - 55.5
 GNN-Ret (K=10) 30.0 14.4 24.4 35.2 19.6 44.0 47.7 32.9 46.3 - - 58.2
-GNN-Ret (w. train) 31.1 17.1 27.3 36.1 21.9 44.4 48.1 33.5 46.9 - - -
+GNN-Ret (w. train) 31.1 17.5 27.3 36.1 21.9 44.4 48.1 33.5 46.9 - - -
 Multi-hop Retrieval
-SelfAsk [23] 28.0 18.3 22.9 37.1 29.6 36.9 51.2 41.5 44.8 - - -
-ITER-RETGEN [27] 30.0 16.7 27.1 32.6 20.0 39.6 45.6 33.3 43.8 - - -
-IRCoT [30] 29.9 12.5 27.5 32.5 18.8 41.0 46.0 30.8 44.6 - - -
-KGP [33] 30.0 15.4 24.0 33.9 18.5 42.1 45.5 31.9 45.2 - - -
+SelfAsk 28.0 18.3 22.9 37.1 29.6 36.9 51.2 41.5 44.8 - - -
+ITER-RETGEN 30.0 16.7 27.1 32.6 20.0 39.6 45.6 33.3 43.8 - - -
+IRCoT 29.9 12.5 27.5 32.5 18.8 41.0 46.0 30.8 44.6 - - -
+KGP 30.0 15.4 24.0 33.9 18.5 42.1 45.5 31.9 45.2 - - -
 RGNN-Ret 32.9 20.8 31.3 43.4 28.3 48.1 59.7 43.3 55.8 - - -
@@
 Table 1: F1/EM/Acc for different QA methods with ChatGPT on four QA datasets. The best and second best scores are highlighted in bold and underline, respectively
+[…]
 4.2 Main Results
 Table 1 summarizes the results for our proposed methods and baselines on four QA datasets using ChatGPT.
-The results show that our proposed GNN-Ret significantly outperforms baselines Dense and Direct in terms of F1, EM, and accuracy on all these four QA datasets with a single retrieval.
-For instance, compared with Dense, GNN-Ret improves EM by 4.4 and accuracy by $4\%$ on the IIRC dataset.
+Results show that GNN-Ret significantly outperforms baselines in terms of F1, EM, and accuracy on all these four QA datasets with a single retrieval.
+For instance, compared with SBERT, GNN-Ret improves EM by 4.4 and accuracy by $4\%$ on the IIRC dataset.
 Additionally, GNN-Ret maintains its superiority on the Quality dataset, which requires a comprehensive understanding of the context of an entire story.
-Notably, GNN-Ret effectively enhances accuracy by $4.8\%$ with $K=10$ on the Quality dataset.
-This improvement can be attributed to the ability of GNN-Ret to enhance the retrieval prioritization of supporting passages through the incorporation of structural information, thereby improving the comprehension of LLMs in articles.
-Surprisingly, our proposed one-hop GNN-Ret even slightly outperforms baselines that perform multi-hop retrieval processes and queries of LLMs in accuracy, which highlights the significant potential of leveraging passage relatedness to enhance the retrieval process for answering multi-hop reasoning questions.
-Furthermore, the results in Table 1 demonstrate that our proposed RGNN-Ret achieves state-of-the-art performance in terms of F1, EM, and accuracy on the three multi-hop reasoning datasets.
-RGNN-Ret outperforms the best baseline, KGP, by more than $10\%$ in accuracy on 2WikiMQA dataset.
+This can be attributed to the ability of GNN-Ret to enhance the retrieval prioritization of supporting passages through the incorporation of structural information.
+Surprisingly, our proposed one-hop GNN-Ret even slightly outperforms baselines with multi-hop retrieval processes and queries of LLMs in accuracy, highlighting significant potentials of leveraging passage relatedness to enhance the retrieval process for answering multi-hop reasoning questions.
+Furthermore, results in Table 1 demonstrate that our proposed RGNN-Ret achieves state-of-the-art performance in terms of F1, EM, and accuracy on the three multi-hop reasoning datasets.
+RGNN-Ret outperforms the best baseline, KGP, by more than $10\%$ in accuracy on 2WikiMQA.
 This performance improvement can be attributed to our proposed self-critique technique and RGNN.
-The self-critique technique enables more accurate judgement on the ending of the answering process and the generation of final answers, and the incorporation of RGNN further enhances the retrieval process for each step of answering, which is essential for generating subquestions and subanswers.
+The self-critique technique enables more accurate judgement on the ending of the answering process and the generation of final answers.
+The incorporation of RGNN further enhances retrieval for each step of answering.
 4.3 Additional Analysis
-Selection of $K$ and $O$.
-In order to effectively train the GNN and RGNN models, we conduct ablation studies to select hyperparameters $K$ and $O$, which determine the number of relevant nodes to be sampled and the size of the competitive node set, respectively.
-To assess the impact of different values of $K$ and $O$, we evaluate the accuracy of GNN-Ret and RGNN-Ret on the MuSiQue and IIRC datasets with varying $K$ and $O$.
-The average accuracy across these two datasets is represented by the green points in Fig.
-4.
-The results demonstrate that both GNN-Ret and RGNN-Ret consistently achieve stable accuracy performance across different settings of $K$ and $O$.
-Upon analyzing the results, we observe that GNN-Ret achieves the highest accuracy when $K=5$ and $O=25$, while RGNN-Ret performs best with $K=5$ and $O=10$.
-GNN-Ret adopts a larger value of $O$ since it considers the whole label set during training, while RGNN-Ret only considers the labels for the current step and subsequent steps.
-Therefore, we select these settings of $K$ and $O$ to train the GNN and RGNN models for experiments reported in Table 1.
-(a) GNN-Ret
-(b) RGNN-Ret
-Figure 4: Accuracy of GNN-Ret and RGNN-Ret with various $K$ and $O$ in Musique and IIRC datasets. The average accuracy of two datasets are displayed in green points.
 Effectiveness of components in GNN-Ret.
@@
 Effectiveness of components in RGNN-Ret.
-We conduct ablation studies to evaluate the effectiveness of each component in RGNN-Ret, and present results in Table 2.
-We first remove the RGNN and solely performs dense retrieval for our proposed self-critique technique.
-The results in Table 2 show that it results in an accuracy reduction of $2.1\%$ and $2.7\%$ on MuSiQue and 2WikiMQA datasets, respectively.
-Next, we incorporate the GNN for the retrieval process of each subquestion, i.e., Self-critique + GNN.
-This modification improves accuracy on both MuSiQue and 2WikiMQA datasets compared to the settings with dense retrieval, indicating the effectiveness of the GNN in enhancing the retrieval coverage for each subquestion individually.
-By combining self-critique and RGNN, our proposed method RGNN-Ret achieves accuracy up to $31.3\%$ and $55.6\%$ on MuSiQue and 2WikiMQA datasets, respectively, which demonstrates the effectiveness of RGNN-Ret in enhancing the retrieval by interconnecting with the graph of passages from previous steps.
+We conduct ablation studies to evaluate the effectiveness of each component in RGNN-Ret (i.e., recurrent, GNN, and RGNN).
+Results in Table 2 show that both the recurrent part (parameter $\beta$) and GNN improve accuracy compared with setting with SBERT on MuSiQue and 2WikiMQA datasets.
+By combining the recurrent part and GNN, our proposed method RGNN-Ret further improves accuracy up to $31.3\%$ and $55.6\%$ on MuSiQue and 2WikiMQA datasets, respectively, which demonstrates the effectiveness of RGNN-Ret in enhancing the retrieval by interconnecting with the GoPs from previous steps.
 This interconnection enhances the retrieval of supporting passages across multiple steps, leading to more accurate answers in the multi-hop reasoning process.
-Method MuSiQue 2WikiMQA
-Dense 23.8 42.3
-GNN-Ret (mean) 26.3 44.6
-GNN-Ret 27.3 46.9
-Self-critique + dense 29.2 52.9
-Self-critique + GNN 30.8 55.2
-RGNN-Ret 31.3 55.6
+Methods MuSiQue 2WikiMQA
+SBERT 23.8 42.3
+GNN-Ret (mean) 26.3 (+2.5) 44.6 (+1.7)
+GNN-Ret 27.3 (+3.5) 46.9 (+4.6)
+Self-critique + SBERT 27.9 52.9
+Self-critique + recurrent 28.1 (+0.2) 54.0 (+1.1)
+Self-critique + GNN 30.8 (+2.9) 55.2 (+2.3)
+RGNN-Ret 31.3 (+3.4) 55.6 (+2.7)
 RGNN-Ret ($Y$) 27.1 53.3
-RGNN-Ret ($Y^{t}$) 30.2 54.6
-RGNN-Ret 31.3 55.6
-Table 2: Ablation studies of components of GNN and RGNN in accuracy with ChatGPT on MuSiQue and 2WikiMQA datasets.
+RGNN-Ret ($Y^{t}$) 30.2 (+3.1) 54.6 (+1.3)
+RGNN-Ret ($Y^{t+}$) 31.3 (+4.2) 55.6 (+2.3)
+Table 2: Ablation studies of components of GNN-Ret and RGNN-Ret in accuracy with ChatGPT on MuSiQue and 2WikiMQA.
 We also explore different settings for the target node sets during the training of RGNN and present results in Table 2.
@@
 This comprehensive approach further improves the accuracy to $31.3\%$ and $55.6\%$ on MuSiQue and 2WikiMQA datasets, respectively, which verifies the effectiveness of RGNN-Ret in enhancing the retrieval process for subsequent subquestions.
-Statistics of retrieval accuracy.
-We explore the EM between the retrieved passages and the ground-truth passages and collect the exact-match number of test samples with varying numbers of supporting passages required in questions on MuSiQue and 2WikiMQA datasets.
-We do not evaluate the statistics of retrieval accuracy for RGNN-Ret without the specific ground-truth supporting passages for each subquestion.
-The results in Fig.
-5 show that our proposed GNN-Ret achieves higher retrieval accuracy compared with Dense on MuSiQue and 2WikiMQA datasets.
-For instance, GNN-Ret outperforms Dense by 29 and 43 exact-match test samples for the questions that require 2 supporting passages.
-This demonstrates that the GNN indeed improves retrieval coverage of supporting passages and the retrieval prioritization of supporting passages for subsequent reasoning.
-We display some case studies to demonstrate advantages of GNN-Ret and RGNN-Ret in Appendix C.
-(a) MuSiQue
-(b) 2WikiMQA
-Figure 5: Exact-match number of test samples with varying numbers of supporting passages required for QA on MuSiQue and 2WikiMQA datasets.
+Effectiveness of GoPs.
+We explore the effectiveness of components (structural information (SI) and shared keywords (SK)) for graph construction.
+Results in Table 3 show that both structural information and shared keywords are able to improve accuracy for the baseline without GoPs.
+Notably, in the context-specific Quality story dataset, which demands a comprehensive grasp of the entire narrative for accurate retrieval, the graph that incorporates structural information markedly outperforms the setting that relies solely on shared keywords.
+While for 2WikiMQA that requires knowledge of multiple documents, the graph with shared keywords achieves a better accuracy compared with that with structural information.
+By combining them into graph construction, we achieve the best accuracy in these two datasets.
+Methods 2WikiMQA Quality
+SBERT (w.o. graph) 42.3 53.4
+graph w. SI 42.7 (+0.4) 55.5 (+2.1)
+graph w. SK 45.4 (+3.1) 54.8 (+1.4)
+graph w. SI and SK 46.9 (+4.6) 58.2 (+4.8)
+Table 3: Ablation studies of components on graph construction in accuracy with ChatGPT on 2WikiMQA and Quality. SI and SK represent structural information and shared keywords, respectively.
+Performance of GNN-Ret with open-sourced LMs.
+Our proposed GNN-Ret utilizes the relatedness of passages to enhance retrieval with the meticulously designed GNNs, which is generalized and robust to all the LM architectures.
+To substantiate this, we supplement the experiments for GNN-Ret and SBERT with two well-behaved open-sourced LMs (Qwen/Qwen2-7B-Instruct Yang et al.
+(2024) and google/gemma-2-9b-it Team (2024)) from Open LLM Leaderboard Fourrier et al.
+(2024).
+Results in Table 4 show that GNN-Ret consistently outperforms SBERT across varying open-sourced LMs, demonstrating the generalization of GNN-Ret in enhancing retrieval coverage with varying LMs.
+MuSiQue 2WikiMQA
+Qwen2-7B-Instruct F1 EM Acc F1 EM Acc
+SBERT 27.8 16.9 23.3 43.3 34.0 40.8
+GNN-Ret 32.3 20.8 26.7 50.5 37.3 45.8
+gemma-2-9b-it F1 EM Acc F1 EM Acc
+SBERT 32.8 18.9 26.9 46.4 28.8 43.8
+GNN-Ret 34.6 20.8 29.0 51.6 30.8 45.8
+Table 4: F1/EM/Acc of SBERT and GNN-Ret with open-sourced LMs on MuSiQue and 2WikiMQA.
 5 Conclusion
@@
 By extending GNN-Ret to multi-hop reasoning questions, we proposed RGNN-Ret, which enhances the retrieval for subquestions through the interconnection between graphs of passages across steps.
-The experiments clearly demonstrated the superiority of both GNN-Ret and RGNN-Ret over baselines, highlighting the effectiveness of leveraging the relatedness between passages to enhance the retrieval process.
+The experiments clearly demonstrated the superiority of both GNN-Ret and RGNN-Ret over baselines, highlighting the effectiveness of leveraging relatedness between passages to enhance retrieval.
 From these advantages, we believe that the incorporation of graph-based representations and the exploitation of passage relatedness can open up new avenues of research in the field of LLMs in understanding structural documents and answering complex questions.
+6 Limitations
+Costs of graph construction.
+While our proposed GNN-Ret demonstrates impressive accuracy in QA tasks with a single query of LLMs, it still relies on multiple LLM queries to extract keywords from passages and construct the graph of passages.
+Fortunately, there are alternative fine-tuned language models specific for keyword extraction [^2] https://huggingface.co/ml6team/keyphrase-extraction-distilbert-inspec [^3] https://github.com/MaartenGr/KeyBERT.
+These models are more compact and significantly accelerate the graph construction process.
+Moreover, the flexibility of the graph of passages allows for dynamic modifications, such as adding new passage nodes or removing outdated ones.
+By maintaining a domain-specific graph of passages, we can effectively address various questions for this domain without the need for reconstructing the graph.
+Costs of message passing.
+We first analyze the statistics of the graph of passages in experiments and detail results in Table 5.
+In larger datasets, each node typically exhibits a higher number of neighbors, attributed to an increased likelihood of keyword sharing among more nodes.
+We calculate the density of the graph of passages, denoted as $D=\frac{2|\mathcal{E}|}{|\mathcal{V}|(|\mathcal{V}|-1)}$, for each dataset in our experiments, based on the established definition Pieterse and Black (2005), where $\mathcal{E}$ and $\mathcal{V}$ represent the edges and vertices of the graph, respectively.
+The densities of all graphs are significantly less than 1, underscoring the substantial sparsity observed in these graphs across our experimental datasets.
+We adopt relevant nodes sampling (Section 3.2) to allow the relevant node set with only top $K$ smallest semantic distances to pass the information to their neighbors during message passing.
+Assuming an average number of neighbors $\mathcal{V}_{K}$ among these $K$ relevant nodes, the time complexity for message passing can be derived as $O(K|\mathcal{V}_{K}|)$, considering a single-layer GNN.
+With $K$ set to 5 or 10 in our experiments, and given the small empirical average number of edges per node $\frac{|\mathcal{E}|}{|\mathcal{V}|}$ reported in Table 5, the time complexity of message passing is neglectable compared with LLM queries.
+Datasets Number of nodes ($|\mathcal{V}|$) Average number of edges ($\frac{|\mathcal{E}|}{|\mathcal{V}|}$) Density
+2WikiMQA 9815 91.38 0.018
+IIRC 21866 259.22 0.024
+MuSiQue 20071 335.77 0.033
+Quality 1104 26.00 0.047
+Table 5: Statistics of graph of passages in experiments.
+Undirected Graph.
+In this study, our focus lies solely on improving the retrieval process through an undirected graph of passages, with no additional information associated with the edges.
+While we employ a sampling technique to select relevant nodes and utilize the minimum semantic distance as the message in the GNN, we acknowledge that there is still a possibility of irrelevant passages receiving messages, potentially impacting the retrieval performance.
+Consequently, the task of selecting an appropriate path on the graph of passages that aligns with the given question remains an area that is yet to be thoroughly explored.
+In addition, we only consider the structural-related passages that are in the same section.
+However, some questions may require more complex structural information.
+For example, when the question is about comparison of two entities in different sections, more complex structural information is needed to connect these corresponding passages to enhance retrieval.
+Therefore, a more delicate graph of passages for more complex tasks warrant further investigation in future studies.
 Appendix A Training Details
@@
 With the ground-truth set of supporting passages $\mathcal{S}_{Y}$, we first define the average integrated semantic distances of supporting passages $\bar{d}_{Y}^{L}$ and non-target nodes $\bar{d}_{O/Y}^{L}$ as below:
-$\bar{d}_{Y}^{L}=\frac{1}{|\mathcal{S}_{Y}|}\sum_{j\in\mathcal{S}_{Y}}h_{j}^{L},\quad\bar{d}_{O/Y}^{L}=\frac{1}{|\mathcal{S}_{O/Y}|}\sum_{o\in\mathcal{S}_{O/Y}}h_{o}^{L},$ (8)
+$\bar{d}_{Y}^{L}=\frac{1}{|\mathcal{S}_{Y}|}\sum_{j\in\mathcal{S}_{Y}}h_{j}^{L},$ (6)
+$\bar{d}_{O/Y}^{L}=\frac{1}{|\mathcal{S}_{O/Y}|}\sum_{o\in\mathcal{S}_{O/Y}}h_{o}^{L},$ (7)
 where $S_{O}$ is the competitive node set with the top $O$ smallest semantic distances and $S_{O/Y}$ is the non-target node set by removing the target node set $S_{Y}$ from $S_{O}$, respectively.
@@
 $\ell=\max(0,r+\bar{d}_{Y}^{L}-\bar{d}_{O/Y}^{L}).$ (8)
-We update the parameters of GNN using gradient descent, and the gradient of $\ell$ with respect to $\alpha^{l}$ is given by:
-$\frac{\partial\ell}{\partial\alpha^{l}}=\left\{\begin{aligned} 0,\quad\ell<=0\\ f_{\alpha^{l}}(L,\mathcal{S}_{Y}),\quad\ell>0\end{aligned}\right.,$ (10)
-where
-$f_{\alpha^{l}}(L,\mathcal{S}_{Y})$ $=\frac{\sum_{j\in\mathcal{S}_{Y}}(h_{j}^{l-1}-m_{j}^{l})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{j}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{Y}|}$ (11)
-$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}_{O/Y}^{l}}(h_{o}^{l-1}-m_{o}^{l})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{o}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{O/Y}^{l}|}.$
+We update the parameters of GNN using gradient descent, and the backward process is presented as below.
+With only partial nodes accepting messages from neighbors, we first define the received node set at each layer $l$ as follows:
+$\mathcal{S}_{R}^{l}=\left\{\begin{aligned} \mathcal{S}_{O}\cap\mathcal{S}_{K}^{L},\quad l=L\\ \mathcal{S}_{K}^{l},\quad\text{else}\end{aligned}\right..$ (9)
+For each node $i\in\mathcal{S}_{R}^{L}$ at the last layer $L$, the gradient of $\ell$ w.r.t. $h_{i}^{L}$ is:
+$\frac{\partial\ell}{\partial h^{L}_{i}}=\left\{\begin{aligned} 1.0,\quad i\in\mathcal{S}_{Y}\\ -1.0,\quad\text{else}\end{aligned}\right..$ (10)
+For the layer $l<L$, the gradient of $l$ w.r.t. $\alpha^{l}$ is presented as follows:
+$\frac{\partial\ell}{\partial\alpha^{l}}=\sum_{j\in\mathcal{S}_{R}^{l}}\frac{\partial\ell}{\partial h^{l}_{i}}(h_{i}^{l-1}-m_{i}^{l}).$ (11)
+Assuming that node $i\in\mathcal{T}_{k}^{l-1}$ at layer $l-1$ transfers the message to the node set $\mathcal{S}_{K,i}^{l}$ at layer $l$ (one node can transfers messages to multiple neighbors), we update the gradient of $\alpha^{l}$ w.r.t.
+$h^{l-1}_{i}$ at the previous layer $l-1$ as follows:
+$\frac{\partial\ell}{\partial h^{l-1}_{i}}=\sum_{j\in\mathcal{S}_{K,i}^{l}}-\alpha^{l}\frac{\partial\ell}{\partial h_{j}^{l}}+\mathbb{1}(i\in\mathcal{S}_{K,i}^{l})\alpha^{l}\frac{\partial\ell}{\partial h_{j}^{l}}.$ (12)
+By propagating the gradients backward to the first layer, we can obtain the gradients for each layer of $\alpha$.
 A.2 Training of RGNN
@@
 We first define the average semantic distances of supporting passages ${\bar{d}_{Y^{t+}}^{L,t}}$ and non-target nodes $\bar{d}_{O/Y^{t+}}^{L,t}$ at step $t$ as below:
-$\bar{d}_{Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{Y^{t+}}|}\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\hat{h}_{j}^{L,t},\quad\bar{d}_{O/Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{O/Y^{t+}}|}\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\hat{h}_{o}^{L,t},$ (12)
+$\bar{d}_{Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{Y^{t+}}|}\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\hat{h}_{j}^{L,t},$ (13)
+$\bar{d}_{O/Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{O/Y^{t+}}|}\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\hat{h}_{o}^{L,t},$ (14)
 where $\mathcal{S}_{O}^{t}$ is the competitive node set with top $O$ smallest semantic distances at step $t$ and $\mathcal{S}_{O/Y^{t+}}^{t}$ is the non-target node set by removing the target node set $\mathcal{S}_{Y^{t+}}^{t}$ from $\mathcal{S}_{O}^{t}$, respectively.
 We formulate the hinge objective function with threshold $r$ for the RGNN as follows:
-$\ell$ $=\frac{1}{T}\sum_{t\in[T]}\ell^{t}=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (13)
-The gradient of loss $\ell^{t}$ at step $t$ with respect to $\alpha$ is computed by:
-$\frac{\partial\ell^{t}}{\partial\alpha^{l}}=\left\{\begin{aligned} 0,\quad\ell<=0\\ \sum_{\tau=0}^{t}\beta(1-\beta)^{t-\tau}f_{\alpha^{l}}(L,\mathcal{S}^{t}_{Y^{t+}}),\quad\ell>0\end{aligned}\right.,$ (14)
+$\ell$ $=\frac{1}{T}\sum_{t\in[T]}\ell^{t}$
+$=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (15)
+We update the parameters of RGNN using gradient descent, and the backward process is presented as below.
+With only partial nodes accepting messages from neighbors, we first define the recieved node set $\mathcal{S}_{R}^{l,t}$ at layer $l$ at step $t$ as follows:
+$\mathcal{S}_{R}^{l,t}=\left\{\begin{aligned} \mathcal{S}_{O}^{t}\cap\mathcal{S}_{K}^{L,t},\quad l=L\\ \mathcal{S}_{K}^{l,t},\quad\text{else}\end{aligned}\right..$ (16)
+For each node $i\in\mathcal{S}_{R}^{L,t}$ at the last layer $L$, the gradient of $\ell^{t}$ w.r.t. $\hat{h}_{i}^{L,t}$ is:
+$\frac{\partial\ell^{t}}{\partial\hat{h}^{L,t}_{i}}=\left\{\begin{aligned} 1.0,\quad i\in\mathcal{S}_{Y^{t}}^{t}\\ -1.0,\quad\text{else}\end{aligned}\right..$ (17)
+Therefore, the gradient of $\ell$ w.r.t. $\hat{h}_{i}^{L,t}$ is:
+$\frac{\partial\ell}{\partial\hat{h}^{L,t}_{i}}=\sum_{\tau=t}^{T}\frac{\partial\ell^{\tau}}{\partial\hat{h}^{L,t}_{i}}=\frac{\partial\ell^{t}}{\partial\hat{h}^{L,t}_{i}}+\frac{\partial\ell}{\partial\hat{h}^{L,t+1}_{i}}(1-\beta),$ (18)
 where
-$f_{\alpha^{l}}(L,\mathcal{S}^{t}_{Y^{t+}})$ $=\frac{\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}(h_{j}^{l-1,t}-m_{j}^{l,t})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{j}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}^{t}_{Y^{t+}}|}$ (15)
-$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}_{O/Y}^{l}}(h_{o}^{l-1,t}-m_{o}^{l,t})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{o}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{O/Y}^{l}|}.$
-The gradient of loss $\ell^{t}$ at step $t$ with respect to $\beta$ is computed by:
-$\frac{\partial\ell^{t}}{\partial\beta}=\left\{\begin{aligned} 0,\quad\ell<=0\\ f_{\beta}(t,L,\mathcal{S}^{t}_{Y^{t+}}),\quad\ell>0\end{aligned}\right.,$ (16)
-where
-$f_{\beta}(t,L,\mathcal{S}^{t}_{Y^{t+}})$ $=\frac{\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\sum_{\tau=1}^{t}(h_{j}^{L,\tau}-\hat{h}_{j}^{L,\tau-1})(1-\beta)^{t-\tau}}{|\mathcal{S}^{t}_{Y^{t+}}|}$ (17)
-$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\sum_{\tau=1}^{t}(h_{o}^{L,\tau}-\hat{h}_{o}^{L,\tau-1})(1-\beta)^{t-\tau}}{|\mathcal{S}_{O/Y^{t+}}^{t}|}.$
+$\frac{\partial\ell}{\partial\hat{h}^{L,T}_{i}}=\frac{\partial\ell^{T}}{\partial\hat{h}^{L,T}_{i}}.$ (19)
+Each node $j\in\mathcal{V}$ at step $t$ transfers the message to the same node at the next step through $\beta$. Therefore, the gradient of loss $\ell$ with respect to $\beta$ is computed by:
+$\frac{\partial\ell}{\partial\beta}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{|\mathcal{V}|}\sum_{j\in\mathcal{V}}\frac{\partial\ell}{\partial\hat{h}^{L,t}_{j}}(h_{j}^{L,t}-\hat{h}_{j}^{L,t-1}).$ (20)
+Recall that $\hat{h}^{L,t}_{i}=\beta{h}^{L,t}_{i}+(1-\beta){h}^{L,t-1}_{i}$ for each node $i\in\mathcal{V}$. The gradient of loss $\ell$ with respect to ${h}^{L,t}_{i}$ is computed by:
+$\frac{\partial\ell}{\partial{h}^{L,t}_{i}}=\beta\frac{\partial\ell}{\partial\hat{h}^{L,t}_{i}}.$ (21)
+With $\frac{\partial\ell}{\partial{h}^{L,t}_{i}}$, we propagate the gradient backward to the GNN and update the parameter $\alpha^{l}$ layer by layer using (11) and (12).
 Appendix B Additional Experimental Details
+Datasets processing.
+We measure all the methods on four different QA datasets, including multi-hop Wikipedia reasoning datasets: (1) MuSiQue Trivedi et al.
+(2022), (2) IIRC Ferguson et al.
+(2020), (3) 2WikiMQA Ho et al.
+(2020), and a single-hop multi-choice QA dataset: (4) Quality Pang et al.
+(2021).
+For each multi-hop dataset, we randomly sample 500 questions from the development set, which provides the ground-truth supporting passages.
+We use 20 of them to train the GNN and select the rest as test data.
+We use all the Wikipedia documents of these 500 questions as the retrieval corpus, which is significantly larger than that provided by datasets [^4] Prior works often use 10–20 paragraphs or documents as the retrieval corpus for each question Shao et al.
+(2023); Wang et al.
+(2024)..
+For Quality, we randomly sample 30 articles with overall corresponding 561 questions and use these articles as the retrieval corpus.
+For each document or article, we split it into multiple passages with a maximum token length of 200 Sarthi et al.
+(2024).
 Graph of passages construction.
@@
 The passages share the same links of keywords are considered as keyword-related and connected together in the graph.
+(a) GNN-Ret
+(b) RGNN-Ret
+Figure 4: Accuracy of GNN-Ret and RGNN-Ret with various $K$ and $O$ in Musique and IIRC datasets. The average accuracy of two datasets are displayed in green points.
 Implemental details of training.
-We initialize $\alpha^{1}=1.0$ and $\beta=1.0$, and use gradient descent to update the GNN and RGNN.
+We initialize $\alpha=0.1$ for GNN-Ret and $\alpha^{1}=1.0$ and $\beta=1.0$ for RGNN-Ret, respectively.
+We then use gradient descent to update the GNN and RGNN.
 For each iteration, we compute the average gradient of all the training samples and update the parameters with a learning rate of 1.0.
@@
 Since training the RGNN requires ground-truth subquestions, we manually generate subquestions for 5 questions sampled from the preset training data.
-The training samples of MuSiQue, IIRC, and 2WikiMQA datasets for the RGNN are displayed in Appendix G.
-We also display the instructions and prompt templates of all the methods in Appendix F.
-Appendix C Qualitative Results
+The training samples of MuSiQue, IIRC, and 2WikiMQA datasets for the RGNN are displayed in Appendix J.
+We also display the instructions and prompt templates of all the methods in Appendix I.
+Appendix C Selection of $K$ and $O$.
+In order to effectively train the GNN and RGNN models, we conduct grid search for hyperparameters $K$ and $O$, which determine the number of relevant nodes to be sampled and the size of the competitive node set, respectively.
+To assess the impact of different values of $K$ and $O$, we evaluate the accuracy of GNN-Ret and RGNN-Ret on the MuSiQue and IIRC datasets with varying $K$ and $O$.
+The average accuracy across these two datasets is represented by the green points in Fig.
+4.
+Results demonstrate that both GNN-Ret and RGNN-Ret consistently achieve stable accuracy performance across different settings of $K$ and $O$.
+Upon analyzing the results, we observe that GNN-Ret achieves the highest accuracy when $K=5$ and $O=25$, while RGNN-Ret performs best with $K=5$ and $O=10$.
+GNN-Ret adopts a larger value of $O$ since it considers the whole label set during training, while RGNN-Ret only considers the labels for the current step and subsequent steps.
+Therefore, we select these settings of $K$ and $O$ to train the GNN and RGNN models for experiments reported in Table 1.
+(a) MuSiQue
+(b) 2WikiMQA
+Figure 5: Exact-match number of test samples with varying numbers of supporting passages required for QA on MuSiQue and 2WikiMQA datasets.
+Appendix D Statistics of retrieval accuracy.
+We explore the EM between the retrieved passages and the ground-truth passages and collect the exact-match number of test samples with varying numbers of supporting passages required in questions on MuSiQue and 2WikiMQA datasets.
+We do not evaluate the statistics of retrieval accuracy for RGNN-Ret without the specific ground-truth supporting passages for each subquestion.
+Results in Fig.
+5 show that our proposed GNN-Ret achieves higher retrieval accuracy compared with SBERT on MuSiQue and 2WikiMQA datasets.
+For instance, GNN-Ret outperforms SBERT by 29 and 43 exact-match test samples for the questions that require 2 supporting passages.
+This demonstrates that the GNN indeed improves retrieval coverage of supporting passages and the retrieval prioritization of supporting passages for subsequent reasoning.
+Token number for retrieval 1k 3k 5k 10k 20k
+SBERT 18.3 23.8 22.5 22.5 27.9
+GNN-Ret 20.4 25.8 25.8 26.7 31.9
+Table 6: Accuracy of GNN-Ret and SBERT with varying token numbers for retrieval using Qwen/Qwen2-7B-Instruct on MuSiQue dataset.
+Appendix E Ablation Studies of Token Number for Retrieval
+We conduct experiments with varying numbers of tokens for retrieval on SBERT and GNN-Ret with the long-context model Qwen/Qwen2-7B-Instruct Yang et al.
+(2024), which can handle more than 100k tokens.
+Results in Table 6 show that our proposed GNN-Ret consistently outperforms SBERT with varying numbers of tokens for retrieval, demonstrating the robustness of our method for long contexts.
+GNN-Ret Learned values of $\alpha$ MuSiQue Learned values of $\alpha$ 2WikiMQA
+F1 EM Acc F1 EM Acc
+1 layer $\alpha=[0.326]$ 31.1 17.5 27.3 $\alpha=[0.277]$ 48.1 33.5 46.9
+2 layers $\alpha=[0.564,0.821]$ 31.5 16.3 27.3 $\alpha=[0.582,0.578]$ 47.0 33.3 46.5
+3 layers $\alpha=[0.506,0.511,0.944]$ 30.8 16.3 26.7 $\alpha=[0.504,0.597,0.802]$ 48.6 34.4 47.5
+Table 7: F1, EM, and Accuracy of GNN-Ret with varying layers on MuSiQue and 2WikiMQA datasets.
+Appendix F Extending GNN-Ret to Multiple Layers
+Our proposed GNNs are designed to be extensible across multiple layers.
+To explore this capability, we conducted additional experiments by increasing the number of GNN layers.
+The results, along with the final learned values of $\alpha$, are presented in Table 7.
+The results indicate that GNN-Ret maintains consistent performance across different layer configurations.
+This observation could be attributed to the fact that many questions in datasets require only two-hop supporting passages, suggesting that a single-layer GNN is adequately equipped to address these queries.
+Method MuSiQue 2WikiMQA
+IRCoT 27.5 44.6
+IRCoT + RGNN-Ret 30.4 (+2.9) 47.5 (+2.9)
+ITER-RETGEN 27.1 43.8
+ITER-RETGEN + RGNN-Ret 29.8 (+2.7) 47.5 (+3.7)
+Table 8: Accuracy of IRCoT (w. RGNN-Ret) and ITER-RETGEN (w. RGNN-Ret) on MuSiQue and 2WikiMQA datasets.
+Appendix G Effectiveness of RGNN-Ret for other multi-hop answering methods
+To further demonstrate the effectiveness of RGNN-Ret, we supplement experiments by employing it with other multi-hop baselines IRCoT and ITER-RETGEN.
+These methods utilize the generated reasons for retrieval, followed by the generation of the next-step reason or the final answer.
+Results in Table 8 show that RGNN-Ret consistently and significantly improves the accuracy performance for other multi-hop baselines, which demonstrates the effectiveness and adaptability of RGNN-Ret in enhancing the retrieval performance for multi-hop answering processes.
+Appendix H Qualitative Results
 We analyze the qualitative results in experiments and demonstrate the effectiveness of GNN-Ret and RGNN-Ret in improving accuracy for QA in this section.
 GNN-Ret improves the retrieval coverage of supporting passages.
-We compare the retrieval process between Dense and GNN-Ret and display the results on Table 3.
-GNN-Ret can retrieve all of the supporting passages while Dense fails in both of these two cases.
-For the first question, Dense can retrieve the knowledge that ‘the director of file Hotel By The Hour is Rolf Olsen.’.
+We compare the retrieval process between SBERT and GNN-Ret and display the results on Table 9.
+GNN-Ret can retrieve all of the supporting passages while SBERT fails in both of these two cases.
+For the first question, SBERT can retrieve the knowledge that ‘the director of file Hotel By The Hour is Rolf Olsen.’.
 Since it considers passages in isolation during retrieval, the supporting passage for the inquiry information (‘which country is Rolf Olsen from’) cannot be retrieved with a poor semantic distance.
@@
 Question: Which country the director of film Hotel By The Hour is from?
-Ground-truth answer: Austria Dense Retrieved passages: (1) Hotel by the Hour (German title:) is a 1970 West German crime film directed by Rolf Olsen and starring Curd Jürgens, Andrea Rau and Corny Collins.
+Ground-truth answer: Austria SBERT Retrieved passages: (1) Hotel by the Hour (German title:) is a 1970 West German crime film directed by Rolf Olsen and starring Curd Jürgens, Andrea Rau and Corny Collins.
 (2) … Final answer: Rolf Olsen is from Germany.
@@
 Question: What is the date of death of the performer of song Goodbye Pork Pie Hat?
-Ground-truth answer: January 5, 1979 Dense Retrieved passages: (1) ‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album.
+Ground-truth answer: January 5, 1979 SBERT Retrieved passages: (1) ‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album.
 (2) … Final Answer: February 17, 2017 GNN-Ret Retrieved passages: (1)‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album.
@@
 However, SelfAsk cannot retrieve the knowledge about ‘other work of William Steig’, which locates in another passage about the book ‘Doctor De Soto’.
-Using dense retrieval fails to retrieve this passage for QA.
+Using SBERT for retrieval fails to retrieve this passage for QA.
 In contrast, RGNN-Ret retrieve this passage since it enhances the semantic distance by integrating with the small semantic distances from the previous step.
@@
 This qualitative example demonstrates that the RGNN can enhance the retrieval over steps for subquestions and thus improve the accuracy for QA.
-Appendix D Limitations and Discussions
-Undirected Graph.
-In this study, our focus lies solely on improving the retrieval process through an undirected graph of passages, with no additional information associated with the edges.
-While we employ a sampling technique to select relevant nodes and utilize the minimum semantic distance as the message in the GNN, we acknowledge that there is still a possibility of irrelevant passages receiving messages, potentially impacting the retrieval performance.
-Consequently, the task of selecting an appropriate path on the graph of passages that aligns with the given question remains an area that is yet to be thoroughly explored.
-In addition, we only consider the structural-related passages that are in the same section.
-However, some questions may require more complex structural information.
-For example, when the question is about comparison of two entities in different sections, more complex structural information is needed to connect these corresponding passages to enhance retrieval.
-Therefore, a more delicate graph of passages for more complex tasks warrant further investigation in future studies.
-Costs of graph construction.
-While our proposed GNN-Ret demonstrates impressive accuracy in QA tasks with a single query of LLMs, it still relies on multiple LLM queries to extract keywords from passages and construct the graph of passages.
-Fortunately, there are alternative fine-tuned language models specific for keyword extraction [^3] https://huggingface.co/ml6team/keyphrase-extraction-distilbert-inspec [^4] https://github.com/MaartenGr/KeyBERT.
-These models are more compact and significantly accelerate the graph construction process.
-Moreover, the flexibility of the graph of passages allows for dynamic modifications, such as adding new passage nodes or removing outdated ones.
-By maintaining a domain-specific graph of passages, we can effectively address various questions for this domain without the need for reconstructing the graph.
-Appendix E Ethics Statement
-We hold a firm conviction that our proposed GNN-Ret and RGNN-Ret have the potential to yield substantial positive outcomes in real-world scenarios associated with retrieval-augmented generation (RAG).
-GNN-Ret and RGNN-Ret significantly enhanced the retrieval coverage of supporting passages, thereby significantly improving the explainability and reliability of LLMs.
-Given the widespread application of LLMs, we anticipate that the implementation of our proposed retrieval techniques will extend their benefits to an increasing number of individuals, thereby catering to a broader range of needs.
-Contrasting the numerous potential benefits of utilizing LLMs, it is crucial to acknowledge the existence of certain negative impacts, such as the generation of hallucinatory outputs and the associated costs.
-Addressing the issue of hallucinations in LLMs requires ongoing efforts to improve retrieval accuracy and provide comprehensive supporting passages to LLMs.
-By ensuring that LLMs have access to a wider range of relevant information, the occurrence of hallucinatory responses can be mitigated.
-Additionally, cost reduction strategies can be implemented by substituting specific steps of GNN-Ret and RGNN-Ret, e.g., keyword extraction.
-The pretrained small LMs could be an effective way to reduce the costs for applications.
-Appendix F Prompts for Experiments
+Appendix I Prompts for Experiments
 We display prompt templates of all the methods in this section.
-F.1 Prompt template for Direct.
+I.1 Prompt template for Direct.
 The prompt template for Direct is shown as follows:
-Prompt template for Direct Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
-<Question>
+Prompt template for Direct Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words. <Question>
 where <Question> indicates the question.
-F.2 Prompt templates for Dense, IRCoT, and KGP
-Dense, IRCoT, and KGP employ different retrieval methods but the same prompt templates for QA. The prompt template is shown as follows:
-Prompt template for Dense / IRCoT / KGP Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
-Context: <Context>
-Question: <Question>
+I.2 Prompt templates for dense retrievers
+The prompt template for retrievers (bm25, DPR, and SBERT) is shown as follows:
+Prompt template for retrievers Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words. Context: <Context> Question: <Question>
 where <Context> indicates the retrieved passages.
 We repeat the question before the length retrieved passages when answering the questions on IIRC and 2WikiMQA datasets for better performance.
-F.3 Prompt template for SelfAsk
+I.3 Prompt templates for IRCoT and KGP
+IRCoT and KGP employ different retrieval methods but the same prompt templates for QA. The prompt template is shown as follows:
+Prompt template for IRCoT / KGP Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words. Context: <Context> Question: <Question>
+where <Context> indicates the retrieved passages.
+We repeat the question before the length retrieved passages when answering the questions on IIRC and 2WikiMQA datasets for better performance.
+I.4 Prompt template for SelfAsk
 SelfAsk prompts LLMs to generate the follow-up question and answers it with retrieved passages until generating the final answer.
@@
 If no follow up questions are necessary, answer the question in five words directly in format: So the final answer is: xxx.
-Here are the examples of how to answer the questions:
-<Examples>
-Context (Optional): <Context>
-Question:
-<Question>
+Here are the examples of how to answer the questions: <Examples> Context (Optional): <Context> Question: <Question>
 We repeat the question before the length retrieved passages when answering the follow-up questions on IIRC and 2WikiMQA datasets for better performance.
-F.4 Prompt template for ITER-RETGEN
+I.5 Prompt template for ITER-RETGEN
 ITER-RETGEN iteratively answers questions with retrieved passages and uses the generated answer for the next-step retrieval, which continues until the generation of final answer.
 The prompt template is shown as follows:
-Prompt template for ITER-RETGEN Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
-Here are the examples of how to answer the questions:
-<Examples>
-Context: <Context>
-Question:
-<Question>
-Let’s think step by step.
-F.5 Prompt template for RGNN-Ret
+Prompt template for ITER-RETGEN Instruction: Given the following question, create a final answer to the question.
+Please answer less than 6 words.
+Here are the examples of how to answer the questions: <Examples> Context: <Context> Question: <Question> Let’s think step by step.
+I.6 Prompt template for RGNN-Ret
 RGNN-Ret iteratively generates the next-step subquestion, answers the subquestion, and performs self-critique until the generation of final answer.
@@
 Prompt template for generating next-step subquestions Instruction: Your goal is to ask the next step question logically.
-Here are the examples of how to answer the questions:
-<Examples>
-Question:
-<Question>
+Here are the examples of how to answer the questions: <Examples> Question: <Question>
 Prompt template for generating subanswers Instruction: Your goal is to answer the next step question.
 I will provide you some wikipedia snippets, and you need to answer the next step question by considering the wikipedia snippets.
-Context:
-<Context>
-Question:
-<Question>
+Context: <Context> Question: <Question>
 We repeat the question before the length retrieved passages when answering the next-step subanswer on IIRC and 2WikiMQA datasets for better performance.
@@
 If not, output ’Critique: no’.
-Question:
-<Question>
-"Analyze if the intermediate answers is enough to generate the final answer to the question step by step logically. Then, if it is enough, output ’Critique: yes’. If not, output ’Critique: no’.
+Question: <Question> "Analyze if the intermediate answers is enough to generate the final answer to the question step by step logically.
+Then, if it is enough, output ’Critique: yes’.
+If not, output ’Critique: no’.
 The demonstrations of RGNN-Ret for asking the next-step subquestion on MuSiQue, IIRC, and 2WikiMQA datasets are shown as follows:
@@
 Next step: What is the home stadium of Charlton Athletic?
-Appendix G Training data for RGNN
+Appendix J Training data for RGNN
 We manually generate subquestions for 5 questions sampled from the preset training data for MuSiQue, IIRC, and 2WikiMQA datasets. The training samples are shown as follows:
-Training data of MuSiQue for RGNN 1. Question: Why did Roncali leave the place of death of the creator of Malchiostro Annunciation?
-Subquestions: Who is the creator of Malchiostro Annunciation? Where did Titian die? Why did Roncali leave Venice?
-2. Question: Where did who argued that the country of citizenship of Victor Denisov had itself beome an imperialist power declare that he would intervene in the Korean conflict?
+Training data of MuSiQue for RGNN 1.
+Question: Why did Roncali leave the place of death of the creator of Malchiostro Annunciation?
+Subquestions: Who is the creator of Malchiostro Annunciation?
+Where did Titian die?
+Why did Roncali leave Venice?
+2.
+Question: Where did who argued that the country of citizenship of Victor Denisov had itself beome an imperialist power declare that he would intervene in the Korean conflict?
 Subquestions: What is the country of citizenship of Victor Denisov?
@@
 Where did Mao Zedong declare that he would intervene in the Korean conflict?
-3. Question: What military overran much of Erich Zakowski’s place of birth?
-Subquestions: What is the place of birth of Erich Zakowski? What military overran East Prussia?
-4. Question: How many people were in British Colonies where does the london broil cut come from ?
-Subquestions: Where does the london broil cut come from? How many people were in North American?
-5. Question: When was the country established that lies immediately north of the Persian Gulf and the region where the country containing Urim is located?
+3.
+Question: What military overran much of Erich Zakowski’s place of birth?
+Subquestions: What is the place of birth of Erich Zakowski?
+What military overran East Prussia?
+4.
+Question: How many people were in British Colonies where does the london broil cut come from ?
+Subquestions: Where does the london broil cut come from?
+How many people were in North American?
+5.
+Question: When was the country established that lies immediately north of the Persian Gulf and the region where the country containing Urim is located?
 Subquestions: What is the region containing Urim?
@@
 When was Iran established?
-Training data of IIRC for RGNN 1. Question: Was the city where Eva was born the capital of its country?
-Subquestions: Where was Eva born? Is Berlin the capital of its country?
-2. Question: In what state did Galambos attend high school?
-Subquestions: What high school did Galambos attend? In what state is Athens High School located?
-3. Question: What was the previous name of the team that Feng started playing with in 1999?
-Subquestions: What team did Feng start playing with in 1999? What was the previous name of Chongqing Longxin?
-4. Question: How many years after the first Marvel Cinematic Universe film came out was Black Panther released?
-Subquestions: When was the first Marvel Cinematic Universe film released? When was Black Panther released? How many years after Iron Man came out was Black Panther released?
-5. Question: Who was the first draft pick the year Damarius Bilbo went undrafted?
-Subquestions: What year did Damarius Bilbo go undrafted? Who was the first draft pick in 2006?
-Training data of 2WikiMQA for RGNN 1. Question: Which film came out earlier, Subliminal Seduction or Australia Marches With Britain?
-Subquestions: When did Subliminal Seduction come out? When did Australia Marches With Britain come out?
-2. Question: Who is the father-in-law of Deuteria?
-Subquestions: Who is the husband of Deuteria? Who is the father of Eusebio?
-3. Question: Who is the mother of the composer of film 404 (Film)?
-Subquestions: Who is the composer of film 404 (Film)? Who is the mother of Ilayaraja?
-4. Question: Which country the composer of film Sergeant Hassan is from?
-Subquestions: Who is the composer of film Sergeant Hassan? Which country is Tamer Karaoğlu from?
-5. Question: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality?
-Subquestions: What is the director of Across The Badlands? Where is Fred F. Sears from? What is the director of A Gutter Magdalene? Where is George Melford from?
+Training data of IIRC for RGNN 1.
+Question: Was the city where Eva was born the capital of its country?
+Subquestions: Where was Eva born?
+Is Berlin the capital of its country?
+2.
+Question: In what state did Galambos attend high school?
+Subquestions: What high school did Galambos attend?
+In what state is Athens High School located?
+3.
+Question: What was the previous name of the team that Feng started playing with in 1999?
+Subquestions: What team did Feng start playing with in 1999?
+What was the previous name of Chongqing Longxin?
+4.
+Question: How many years after the first Marvel Cinematic Universe film came out was Black Panther released?
+Subquestions: When was the first Marvel Cinematic Universe film released?
+When was Black Panther released?
+How many years after Iron Man came out was Black Panther released?
+5.
+Question: Who was the first draft pick the year Damarius Bilbo went undrafted?
+Subquestions: What year did Damarius Bilbo go undrafted?
+Who was the first draft pick in 2006?
+Training data of 2WikiMQA for RGNN 1.
+Question: Which film came out earlier, Subliminal Seduction or Australia Marches With Britain?
+Subquestions: When did Subliminal Seduction come out?
+When did Australia Marches With Britain come out?
+2.
+Question: Who is the father-in-law of Deuteria?
+Subquestions: Who is the husband of Deuteria?
+Who is the father of Eusebio?
+3.
+Question: Who is the mother of the composer of film 404 (Film)?
+Subquestions: Who is the composer of film 404 (Film)?
+Who is the mother of Ilayaraja?
+4.
+Question: Which country the composer of film Sergeant Hassan is from?
+Subquestions: Who is the composer of film Sergeant Hassan?
+Which country is Tamer Karaoğlu from?
+5.
+Question: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality?
+Subquestions: What is the director of Across The Badlands?
+Where is Fred F.
+Sears from?
+What is the director of A Gutter Magdalene?
+Where is George Melford from?